### v1

# Impact of positrons on electrical conductivity of hot and dense astrophysical plasma

Abstract
We study the influence of positrons on the outer crusts of neutron stars and the interiors of white dwarfs, introducing them as a novel component in both the composition of matter and in transport processes. We solve a system of coupled Boltzmann kinetic equations for the electron and positron distribution functions in the relaxation-time approximation, taking into account electron–ion, positron–ion, and electron–positron collisions. The relevant scattering matrix elements are calculated from one-plasmon exchange diagrams, with in-medium polarization tensors derived within the hard–thermal–loop effective theory. Numerical results are obtained for matter composed of carbon nuclei. We find that the conductivity rises with temperature, following a power law $\sigma\propto T^{4}$ in the semi-degenerate regime and $\sigma\propto T$ in the nondegenerate regime, due to the intense creation of thermal electron–positron pairs and the resulting collisions among them. These results highlight the importance of including positrons in the transport properties of heated, dense astrophysical plasmas.
I Introduction
Electrical conductivity of matter at subsaturation densities in neutron stars and cores of white dwarfs play a central role in several astrophysical problems, such as the magnetic field evolution 1992ApJ…395..250G ; 1995JApA…16..217B ; 1996ApJ…458..347M ; Ascenzi2024 , dissipation of magnetohydrodynamic waves 1979ApJ…227..995E ; 2014Ap…..57..530S , etc. Conductivity of dense neutron-star matter has been studied extensively in the cold regime relevant to isolated stars ($T\leq 1$ MeV) Flowers1976 ; Yakovlev1980 ; Urpin1980 ; Flowers1981 ; Itoh1983 ; Itoh1984 ; Nandkumar1984MNRAS ; Sedrakian1987 ; Itoh1993 ; Baiko1998 ; Potekhin1999 ; Shternin2006 ; Itoh2008 , see Ref. Schmitt2018 for a recent review. However, formation of hot crustal matter in compact stars is anticipated by various astrophysical scenarios, such as supernova explosions leading to hot protoneutron stars, remnants of binary neutron star mergers, as well as heating of neutron star crusts in accreting binaries. In this regime, the plasma forms a liquid state of correlated ions, while the relativistic electron gas becomes non-degenerate. This transition triggers intense production of electron–positron pairs, introducing a positron component into the plasma that contributes significantly to its conductivity. Finite temperature properties of such plasma are expected to significantly deviate from zero‑temperature (cold, degenerate) results.
The electrical conductivity of a one-component plasma in the hot regime under magnetic fields was investigated in Ref. Harutyunyan2016 . These results were later employed in Ref. Harutyunyan2018 to estimate the impact of Ohmic dissipation and the Hall effect on the dynamics of binary neutron star mergers. In addition, the mean free path of conducting electrons was used to identify the density–temperature regimes in which the magnetohydrodynamic description breaks down.
The present work extends the analysis of Ref. Harutyunyan2016 to a hot plasma state in which a non-negligible population of positrons coexists with electrons. Although positron contributions were neglected in earlier studies, they become critically important in the outer crust at temperatures $T\gtrsim 1\textrm{MeV}=1.16\times 10^{10}$ K. In this regime, electrons and positrons constitute the primary charge carriers, and electrical conduction is governed mainly by electron–ion, positron–ion, and electron–positron scattering processes.
To describe transport, we employ Boltzmann kinetic theory in the relaxation-time approximation to obtain the non-equilibrium distribution functions of electrons and positrons, fully accounting for electron–ion, positron–ion, and electron–positron collisions. In-medium screening effects are incorporated through the structure factor of the one-component classical plasma and the hard-thermal-loop effective-theory treatment of the plasmon self-energy Braaten1990 ; Braaten1992 ; Harutyunyan2016 .
This work is structured as follows. In Sec. II we derive the electrical conductivity for a multicomponent plasma consisting of electrons, positrons and ions, solving the system of linearized coupled Boltzmann equations for electron and positron distribution functions. In Section III, the matrix elements for electron-ion, positron-ion and positron-electron scattering processes and the resulting collision rates are discussed. Section IV presents the phase structure of crustal matter and numerical results for electrical conductivity. Our results are summarized in Sec. V. Appendix A provides the details of the derivation of the electron-positron scattering matrix element, and Appendix B computes the collision rate due to this scattering channel.
We use the natural units with $\hbar=c=k_{B}=k_{e}=1$, $e=\sqrt{\alpha}$, $\alpha=1/137$ and the metric signature (1,1,1,1)(1,-1,-1,-1). Greek and Latin indices are utilized to label 4-space and 3-space tensor quantities, respectively.
II Electrical conductivity from the Boltzmann equation
The kinetics of electrons and positrons is described by a system of coupled Boltzmann equations
$\frac{\partial f^{-}}{\partial t}+\bm{v}\frac{\partial f^{-}}{\partial\bm{r}}-e(\bm{E}+\bm{v}\times\bm{H})\frac{\partial f^{-}}{\partial\bm{p}}=I^{-},$ (1)
$\frac{\partial f^{+}}{\partial t}+\bm{v}\frac{\partial f^{+}}{\partial\bm{r}}+e(\bm{E}+\bm{v}\times\bm{H})\frac{\partial f^{+}}{\partial\bm{p}}=I^{+},$ (2)
where $f^{-}$ and $f^{+}$ are the distribution functions for electrons and positrons, respectively, $\bm{E}$} and $\bm{H}$} are the electric and magnetic fields, $\bm{v}$} is the electron/positron velocity, and $I^{-}$ and $I^{+}$ are the collision integrals, respectively, for electrons and positrons, which include electron-ion and electron-positron scattering for $I^{-}$, and positron-ion and positron-electron scattering for $I^{+}$
$I^{-}=\sum_{i}I_{ei}+I_{ep},$ (3)
$I^{+}=\sum_{i}I_{pi}+I_{pe},$ (4)
where the sums in the first terms go over the type of ions (here we consider the general case of a multi-component plasma, where more than one sort of ions can be present simultaneously). For electron-ion and positron-ion collision integrals, we have
$I_{ei}$ $=$ $-(2\pi)^{4}\sum\limits_{234}|{\cal M}_{12\to 34}^{ei}|^{2}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})$ (5)
$\hskip-28.45274pt\times\Big[f^{-}_{1}g_{2}^{i}(1-f^{-}_{3})(1\pm g_{4}^{i})-f^{-}_{3}g_{4}^{i}(1-f^{-}_{1})(1\pm g_{2}^{i})\Big],$
$I_{pi}$ $=$ $-(2\pi)^{4}\sum\limits_{234}|{\cal M}_{12\to 34}^{pi}|^{2}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})$ (6)
$\hskip-28.45274pt\times\Big[f^{+}_{1}g_{2}^{i}(1-f^{+}_{3})(1\pm g_{4}^{i})-f^{+}_{3}g_{4}^{i}(1-f^{+}_{1})(1\pm g_{2}^{i})\Big],\quad$
where the indices $1$ and $3$ label the incoming and outgoing electrons/positrons, respectively, $g^{i}_{2,4}\equiv g^{i}(p_{2,4})$ are the ion distribution functions before and after the collision, and ${\cal M}_{12\to 34}^{ei,pi}$ is the electron/positron–ion scattering matrix element. The upper signs in Eqs. (5) and (6) correspond to Bose statistics and the lower signs to Fermi statistics for ions. We also use the shorthand notation $\sum\limits_{i}=\int d{\bm{p}}_{i}/(2\pi)^{3}$. Assuming that ions remain in thermal equilibrium, we take $g^{i}(p)$ in the form of the Bose or Fermi distribution function, $g^{i}(p)=\left[e^{\beta(\varepsilon_{p}-\mu_{i})}\mp 1\right]^{-1},$ where $\varepsilon_{p}=p^{2}/M_{i}$, and $M_{i}$ and $\mu_{i}$ are the mass and chemical potential of ions of type $i$, while $\beta=T^{-1}$ is the inverse temperature.
The electron-positron and positron-electron collision integrals can be written in a similar fashion
$I_{ep}$ $=$ $-(2\pi)^{4}\sum\limits_{234}|{\cal M}_{12\to 34}^{ep}|^{2}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})$ (7)
[…]
$I_{pe}$ $=$ $-(2\pi)^{4}\sum\limits_{234}|{\cal M}_{12\to 34}^{pe}|^{2}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})$ (8)
$\hskip-42.67912pt\times\Big[f^{+}_{1}f^{-}_{2}(1-f^{+}_{3})(1-f^{-}_{4})-f^{+}_{3}f^{-}_{4}(1-f^{+}_{1})(1-f^{-}_{2})\Big],$
where ${\cal M}_{12\to 34}^{ep}$ and ${\cal M}_{12\to 34}^{pe}$ are the electron-positron and positron-electron scattering matrix elements, respectively.
To solve the Boltzmann equations, we will consider weak deviations from equilibrium and search the distribution functions in the form
$f^{\pm}=f^{0\pm}+\delta f^{\pm},\qquad\delta f^{\pm}=-\phi^{\pm}\frac{\partial f^{0\pm}}{\partial\epsilon},$ (9)
where $f^{0\pm}({\epsilon})=[e^{\beta(\epsilon\pm\mu_{e})}+1]^{-1}$ are the local Fermi distribution functions for positrons/electrons with $\epsilon=\sqrt{p^{2}+m^{2}}$, $m$ is the electron mass, $\mu$ is the electron chemical potential, $\delta f^{\pm}\ll f^{0\pm}$ are small perturbations, and $\phi^{\pm}$ are unknown functions; additionally, introducing fermionic velocity $\bm{v}=\partial\epsilon/\partial\bm{p}=\bm{p}/\epsilon$, we quote some of the derivatives that will be used below
$\frac{\partial f^{0\pm}}{\partial\bm{p}}$ $=$ $\bm{v}\frac{\partial f^{0\pm}}{\partial\epsilon},\quad\frac{\partial f^{0\pm}}{\partial\epsilon}=-\beta f^{0\pm}(1-f^{0\pm}).$ (10)
We assume plasma is spatially uniform, time-independent and external fields are uniform and constant, i.e., there is no spatial variation of the fermionic distribution function, with only momentum space evolution.
We next linearize the Boltzmann equation by substituting Eq. (9) into Eqs. (1) and (2) and keeping only the terms linear in $\delta f^{\pm}$ and in the electric field $\bm{E}$}, which implies that in the terms proportional to $\bm{E}$}, the substitutions $f^{\pm}\to f^{0\pm}$ can be made. We further neglect magnetic fields, which induce anisotropy in the conductivity; for the discussion of the effects of $\bm{H}$} field on transport, see Ref. Harutyunyan2016 . Thus, the linearized Boltzmann equations take the form
$-\frac{\partial f^{0-}}{\partial\epsilon}e\bm{v}\cdot\bm{E}=\sum_{i}I_{ei}+I_{ep},$ (11)
$\frac{\partial f^{0+}}{\partial\epsilon}e\bm{v}\cdot\bm{E}=\sum_{i}I_{pi}+I_{pe},$ (12)
where the linearized collision integrals are given by
[…]
$I_{pi}=-\sum_{pi}(\phi^{+}_{1}-\phi^{+}_{3}),~I_{pe}=-\sum_{pe}(\phi^{+}_{1}+\phi^{-}_{2}-\phi^{+}_{3}-\phi^{-}_{4}),$
where
$\sum_{ei}$ $\equiv$ $(2\pi)^{4}\beta\sum\limits_{234}|{\cal M}_{12\to 34}^{ei}|^{2}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})$ (13)
$\times$ $f^{0-}_{1}(1-f^{0-}_{3})g_{2}^{i}(1\pm g_{4}^{i}),$
$\sum_{pi}$ $\equiv$ $(2\pi)^{4}\beta\sum\limits_{234}|{\cal M}_{12\to 34}^{pi}|^{2}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})$ (14)
$\times$ $f^{0+}_{1}(1-f^{0+}_{3})g_{2}^{i}(1\pm g_{4}^{i}),$
$\sum_{ep}$ $\equiv$ $(2\pi)^{4}\beta\sum\limits_{234}|{\cal M}_{12\to 34}^{ep}|^{2}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})$ (15)
$\times$ $f_{1}^{0-}f_{2}^{0+}(1-f_{3}^{0-})(1-f_{4}^{0+}),$
$\sum_{pe}$ $\equiv$ $(2\pi)^{4}\beta\sum\limits_{234}|{\cal M}_{12\to 34}^{pe}|^{2}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})$ (16)
$\times$ $f_{1}^{0+}f_{2}^{0-}(1-f_{3}^{0+})(1-f_{4}^{0-}).\qquad$
We seek solutions of Eqs. (11) and (12) in the form
$\phi^{\pm}={\pm}e\tau^{\pm}(\bm{v}\cdot\bm{E}),$ (17)
where the relaxation times $\tau^{\pm}$ depend on the particle energy $\epsilon$. Substituting this into the collision integrals yields
$I_{ei}$ $=$ $e\bm{E}\sum_{ei}\Big(\tau^{-}_{1}\bm{v}_{1}-\tau^{-}_{3}\bm{v}_{3}\Big),$ (18)
$I_{pi}$ $=$ $-e\bm{E}\sum_{pi}\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{+}_{3}\bm{v}_{3}\Big),$ (19)
$I_{ep}$ $=$ $e\bm{E}\sum_{ep}\Big(\tau^{-}_{1}\bm{v}_{1}-\tau^{+}_{2}\bm{v}_{2}-\tau^{-}_{3}\bm{v}_{3}+\tau^{+}_{4}\bm{v}_{4}\Big),$ (20)
$I_{pe}$ $=$ $-e\bm{E}\sum_{pe}\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{-}_{2}\bm{v}_{2}-\tau^{+}_{3}\bm{v}_{3}+\tau^{-}_{4}\bm{v}_{4}\Big),\quad$ (21)
and Eqs. (11) and (12) become
$-\frac{\partial f^{0-}}{\partial\epsilon}\bm{v}$ $=$ $\sum_{i}\sum_{ei}\Big(\tau^{-}_{1}\bm{v}_{1}-\tau^{-}_{3}\bm{v}_{3}\Big)$ (22)
$\hskip-28.45274pt+\sum_{ep}\Big(\tau^{-}_{1}\bm{v}_{1}-\tau^{+}_{2}\bm{v}_{2}-\tau^{-}_{3}\bm{v}_{3}+\tau^{+}_{4}\bm{v}_{4}\Big).$
$-\frac{\partial f^{0+}}{\partial\epsilon}\bm{v}$ $=$ $\sum_{i}\sum_{pi}\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{+}_{3}\bm{v}_{3}\Big)$ (23)
$\hskip-28.45274pt+\sum_{pe}\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{-}_{2}\bm{v}_{2}-\tau^{+}_{3}\bm{v}_{3}+\tau^{-}_{4}\bm{v}_{4}\Big).\quad$
We multiply Eqs. (22) and (23) by $\tau_{1}^{-}{\bm{v}}_{1}$ and $\tau_{1}^{+}{\bm{v}}_{1}$, respectively, and integrate over ${\bm{p}}_{1}\equiv{\bm{p}}$, yielding,
$-\sum_{1}\frac{\partial f^{0-}}{\partial\epsilon}\tau^{-}{\bm{v}}^{2}$ $=$ $\frac{1}{2}\sum_{1,i}\sum_{ei}\Big(\tau_{1}^{-}\bm{v}_{1}-\tau_{3}^{-}\bm{v}_{3}\Big)^{2}$ (24)
$+$ $\frac{1}{2}\sum_{1}\sum_{ep}\Big(\tau_{1}^{-}\bm{v}_{1}-\tau^{-}_{3}\bm{v}_{3}\Big)$
$\hskip-31.2982pt\Big(\tau^{-}_{1}\bm{v}_{1}-\tau^{+}_{2}\bm{v}_{2}-\tau^{-}_{3}\bm{v}_{3}+\tau^{+}_{4}\bm{v}_{4}\Big),$
$-\sum_{1}\frac{\partial f^{0+}}{\partial\epsilon}\tau^{+}{\bm{v}}^{2}$ $=$ $\frac{1}{2}\sum_{1,i}\sum_{pi}\Big(\tau_{1}^{+}\bm{v}_{1}-\tau_{3}^{+}\bm{v}_{3}\Big)^{2}$ (25)
$+$ $\frac{1}{2}\sum_{1}\sum_{pe}\Big(\tau_{1}^{+}\bm{v}_{1}-\tau^{+}_{3}\bm{v}_{3}\Big)$
$\hskip-34.14322pt\times\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{-}_{2}\bm{v}_{2}-\tau^{+}_{3}\bm{v}_{3}+\tau^{-}_{4}\bm{v}_{4}\Big),\quad$
where we performed permutations $1\leftrightarrow 3$ and $2\leftrightarrow 4$. To simplify the sums we make a simple ansatz $\tau^{\pm}(\epsilon)\propto\epsilon$, which gives ($\epsilon\equiv\epsilon_{1}$)
$\sum_{e,pi}\Big(\tau^{\pm}_{1}\bm{v}_{1}-\tau^{\pm}_{3}\bm{v}_{3}\Big)^{2}=\left(X^{\pm}\right)^{2}\sum_{e,pi}{\bm{q}}^{2},$ (26)
where $\bm{q}=\bm{p}_{1}-\bm{p}_{3}=\bm{p}_{4}-\bm{p}_{2}$ is the transferred momentum, and $X^{\pm}={\tau^{\pm}}(\epsilon)/{\epsilon}={\rm const}$. Similarly,
$\sum_{ep}\Big(\tau_{1}^{-}\bm{v}_{1}-\tau^{-}_{3}\bm{v}_{3}\Big)\Big(\tau^{-}_{1}\bm{v}_{1}-\tau^{+}_{2}\bm{v}_{2}-\tau^{-}_{3}\bm{v}_{3}+\tau^{+}_{4}\bm{v}_{4}\Big)$
$\hskip 56.9055pt=X^{-}\left(X^{-}+X^{+}\right)\sum_{ep}\bm{q}^{2},$ (27)
$\sum_{pe}\Big(\tau_{1}^{+}\bm{v}_{1}-\tau^{+}_{3}\bm{v}_{3}\Big)\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{-}_{2}\bm{v}_{2}-\tau^{+}_{3}\bm{v}_{3}+\tau^{-}_{4}\bm{v}_{4}\Big)$
$\hskip 56.9055pt=X^{+}\left(X^{-}+X^{+}\right)\sum_{pe}\bm{q}^{2}.$ (28)
Then Eqs. (24) and (25) can be written as
$X^{-}\bigg(\sum_{i}\nu_{ei}+\nu_{ep}\bigg)+X^{+}\nu_{ep}=1,$ (29)
$X^{+}\bigg(\sum_{i}\nu_{pi}+\nu_{pe}\bigg)+X^{-}\nu_{pe}=1,$ (30)
with
$\nu_{ei}$ $=$ $\bigg[-\sum_{1}\frac{\partial f^{0-}}{\partial\epsilon}\frac{p^{2}}{\epsilon}\bigg]^{-1}\frac{1}{2}\sum_{1}\sum_{ei}{\bm{q}}^{2},$ (31)
$\nu_{ep}$ $=$ $\bigg[-\sum_{1}\frac{\partial f^{0-}}{\partial\epsilon}\frac{p^{2}}{\epsilon}\bigg]^{-1}\frac{1}{2}\sum_{1}\sum_{ep}{\bm{q}}^{2}.$ (32)
Similar expressions hold for positons with the replacements $ei\to pi$, $ep\to pe$ and $f^{0-}\to f^{0+}$. These quantities represent appropriately weighted mean collision rates between the corresponding particles. Solving (29) and (30) then yields
$X^{-}$ $=$ $\frac{\sum_{i}\nu_{pi}+\nu_{pe}-\nu_{ep}}{\big(\sum_{i}\nu_{ei}+\nu_{ep}\big)\big(\sum_{i}\nu_{pi}+\nu_{pe}\big)-\nu_{ep}\nu_{pe}},$ (33)
$X^{+}$ $=$ $\frac{\sum_{i}\nu_{ei}+\nu_{ep}-\nu_{pe}}{\big(\sum_{i}\nu_{ei}+\nu_{ep}\big)\big(\sum_{i}\nu_{pi}+\nu_{pe}\big)-\nu_{ep}\nu_{pe}}.\quad$ (34)
In the case where electron-positron collisions are neglected, i.e., $\nu_{ep}=\nu_{pe}=0$, Eqs. (33) and (34) reduce to the well-known expressions
$(X^{-})^{-1}=\sum_{i}\nu_{ei},\qquad(X^{+})^{-1}=\sum_{i}\nu_{pi},$ (35)
where the collision rates on different types of nuclei sum up.
In the regimes where electron-positron collisions become important, the number of electrons and positrons is almost equal, and we find that practically always $\nu_{ep}\simeq\nu_{pe}$, such that $\nu_{ep}-\nu_{pe}\ll\nu_{ei}\simeq\nu_{pi}$. Then Eqs. (33) and (34) can be simplified to
$X^{-}$ $=$ $\frac{\sum_{i}\nu_{pi}}{\big(\sum_{i}\nu_{ei}\big)\big(\sum_{i}\nu_{pi}\big)+\nu_{ep}\sum_{i}(\nu_{ei}+\nu_{pi})},$ (36)
$X^{+}$ $=$ $\frac{\sum_{i}\nu_{ei}}{\big(\sum_{i}\nu_{ei}\big)\big(\sum_{i}\nu_{pi}\big)+\nu_{ep}\sum_{i}(\nu_{ei}+\nu_{pi})}.$ (37)
In the opposite limit of fast electron-positron scattering we have $\nu_{ei}\simeq\nu_{pi}\ll\nu_{ep}\simeq\nu_{pe}$, therefore
$X^{-}=X^{+}=\frac{1}{2\nu_{ep}}=\frac{1}{2\nu_{pe}},$ (38)
which can be derived also directly from Eqs. (29) and (30).
We next write explicit expressions for the mean collision rates by substituting Eqs. (13)–(16) into (31) and (32) (and their analogues for positrons) to obtain
$\nu_{ei}$ $=$ $\frac{(2\pi)^{4}}{2l^{-}}\!\sum\limits_{1234}|{\cal M}_{12\to 34}^{ei}|^{2}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})$ (39)
$\times f^{0-}_{1}(1-f^{0-}_{3})g_{2}^{i}(1\pm g_{4}^{i}){\bm{q}}^{2},$
$\nu_{pi}$ $=$ $\frac{(2\pi)^{4}}{2l^{+}}\!\sum\limits_{1234}|{\cal M}_{12\to 34}^{pi}|^{2}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})$ (40)
$\times f^{0+}_{1}(1-f^{0+}_{3})g_{2}^{i}(1\pm g_{4}^{i}){\bm{q}}^{2},$
$\nu_{ep}$ $=$ $\frac{(2\pi)^{4}}{2l^{-}}\sum\limits_{1234}|{\cal M}_{12\to 34}^{ep}|^{2}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})$ (41)
$\times f_{1}^{0-}f_{2}^{0+}(1-f_{3}^{0-})(1-f_{4}^{0+}){\bm{q}}^{2},$
$\nu_{pe}$ $=$ $\frac{(2\pi)^{4}}{2l^{+}}\sum\limits_{1234}|{\cal M}_{12\to 34}^{pe}|^{2}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})$ (42)
$\times f_{1}^{0+}f_{2}^{0-}(1-f_{3}^{0+})(1-f_{4}^{0-}){\bm{q}}^{2},$
with
$l^{\pm}\equiv\frac{1}{2\pi^{2}}\int_{m}^{\infty}\!d\epsilon\,p^{3}f^{0\pm}\left(1-f^{0\pm}\right),$ (43)
where we used Eq. (10).
We can now calculate the contributions of electrons and positrons to the electric current:
$j^{-}_{k}$ $=$ $-\int\!\frac{2d\bm{p}}{(2\pi)^{3}}ev_{k}\,\delta f^{-}=\sigma^{-}_{kj}E_{j},$ (44)
$j^{+}_{k}$ $=$ $\int\!\frac{2d\bm{p}}{(2\pi)^{3}}ev_{k}\,\delta f^{+}=\sigma^{+}_{kj}E_{j},$ (45)
where we used Eqs. (9) and (17), and defined the conductivity matrices (the factor 2 accounts for spin degeneracy) as
$\sigma^{\pm}_{kj}=-\int\!\frac{2d\bm{p}}{(2\pi)^{3}}\frac{\partial f^{0\pm}}{\partial\epsilon}e^{2}\tau^{\pm}v_{k}v_{j}=\delta_{kj}\sigma^{\pm},$ (46)
where isotropy of the medium allows us to define the scalar conductivities:
$\sigma^{\pm}=-\frac{e^{2}}{3\pi^{2}}\int_{m}^{\infty}d\epsilon\,\frac{p^{3}}{\epsilon}\frac{\partial f^{0\pm}}{\partial\epsilon}\tau^{\pm}.$ (47)
The total current is then
$\bm{j}=\bm{j}^{+}+\bm{j}^{-}={\sigma}\bm{E},$ (48)
with total conductivity ${\sigma}={\sigma}^{+}+{\sigma}^{-}$. The electrical conductivity is thus fully determined once the relaxation times $\tau^{\pm}$ are known. Employing the ansatz $\tau^{\pm}=X^{\pm}\epsilon$, Eq. (47) becomes
$\sigma^{\pm}=\frac{e^{2}X^{\pm}}{3\pi^{2}T}\!\int_{m}^{\infty}d\epsilon\,{p^{3}}f^{0\pm}\left(1-f^{0\pm}\right)=\frac{2e^{2}}{3T}X^{\pm}l^{\pm}.$ (49)
Figure 1: Diagrams describing the electron-positron scattering and annihilation via exchange of a virtual plasmon. The plasmon self-energy is given by the polarization tensor $\Pi_{\mu\nu}(\omega,{\bm{q}})$ shown by the closed loops.
III Scattering amplitudes and collision rates
In this section, we will discuss the scattering matrix elements which enter the integrands of collision rates (39)–(42), using the standard QED techniques for a thermal medium. In addition, we will also discuss the electron-positron collision rates using the obtained expressions for the matrix elements.
III.1 The electron-positron scattering matrix element
The screened electron–positron scattering matrix element can be expressed as the sum of two contributions, corresponding to proper scattering and annihilation processes (analogous to Bhabha scattering in vacuum); see Fig. 1.
${\cal M}_{12\to 34}^{ep}$ $=$ ${\cal M}_{12\to 34}^{ep,t}-{\cal M}_{12\to 34}^{ep,s},$ (50)
where each of the channels in a thermal medium can be split into longitudinal and transverse parts (see Ref. Harutyunyan2016 and references therein)
${\cal M}_{12\to 34}^{ep,t}=-{\cal M}^{t}_{L}+{\cal M}^{t}_{T},~{\cal M}^{t}_{L}=\frac{J_{0}J^{\prime}_{0}}{t_{0}},~{\cal M}^{t}_{T}=\frac{\bm{J}_{\perp}\bm{J}^{\prime}_{\perp}}{t_{\perp}},$
${\cal M}_{12\to 34}^{ep,s}=-{\cal M}^{s}_{L}+{\cal M}^{s}_{T},~{\cal M}^{s}_{L}=\frac{\tilde{J}_{0}\tilde{J}^{\prime}_{0}}{s_{0}},~{\cal M}^{s}_{T}=\frac{\tilde{\bm{J}}_{\perp}\tilde{\bm{J}}^{\prime}_{\perp}}{s_{\perp}},$
where we introduced the following (Mandelstam) variables
$t_{0}=q^{2}+\Pi_{L},\quad t_{\perp}=q^{2}-\omega^{2}+\Pi_{T},$ (53)
$s_{0}=\tilde{q}^{2}+\tilde{\Pi}_{L},\quad s_{\perp}=\tilde{q}^{2}-\tilde{\omega}^{2}+\tilde{\Pi}_{T},$ (54)
and 4-currents
$J^{\mu}=-e^{*}\bar{u}_{3}\gamma^{\mu}u_{1},\quad J^{\prime\mu}=-e^{*}\bar{v}_{2}\gamma^{\mu}v_{4},$ (55)
$\tilde{J}^{\mu}=-e^{*}\bar{v}_{2}\gamma^{\mu}u_{1},\quad\tilde{J}^{\prime\mu}=-e^{*}\bar{u}_{3}\gamma^{\mu}v_{4},$ (56)
where $e^{*}=\sqrt{4\pi}e$, $(\omega,{\bm{q}})=p_{1}-p_{3}=p_{4}-p_{2}$, $(\tilde{\omega},\tilde{{\bm{q}}})=p_{1}+p_{2}=p_{3}+p_{4}$, $u_{a}={u}^{s_{a}}(p_{a})$, $v_{b}={v}^{s_{b}}(p_{b})$, $\bm{J}_{\perp},\bm{J}^{\prime}_{\perp}$ are the components of these currents transverse to $\bm{q}$, and $\tilde{\bm{J}}_{\perp},\tilde{\bm{J}}^{\prime}_{\perp}$ are those transverse to $\tilde{\bm{q}}$. The screening of the interaction is taken into account in terms of the longitudinal $\Pi_{L}(\omega,q)$ and transverse $\Pi_{T}(\omega,q)$ components of the polarization tensor.
Details of the squaring and spin-averaging of the $ep$ matrix element are given in Appendix A. In this work, we neglect the interference terms between the $t$- and $s$-channels in the electron–positron scattering amplitude. The final results for the $t$- and $s$-channel spin-averaged scattering amplitudes are
$\frac{1}{4}\sum\limits_{\rm spins}|{\cal M}^{ep,t}_{12\to 34}|^{2}$ $=$ $\frac{e^{*4}}{4\epsilon_{1}\epsilon_{2}(\epsilon_{1}-\omega)(\epsilon_{2}+\omega)}$ (57)
$\hskip-56.9055pt\times\Bigg[\frac{C_{0}(\alpha,\theta)}{|t_{0}|^{2}}+\frac{C_{2}(\alpha,\theta)+4C^{2}_{1}(\alpha,\theta)\cos^{2}\varphi}{|t_{\perp}|^{2}}$
$\hskip-56.9055pt-2{\rm Re}\frac{(2\epsilon_{1}-\omega)(2\epsilon_{2}+\omega)C_{1}(\alpha,\theta)\cos\varphi}{(t_{0}t^{*}_{\perp})}\Bigg],$
$\frac{1}{4}\sum\limits_{\rm spins}|{\cal M}^{ep,s}_{12\to 34}|^{2}$ $=$ $\frac{e^{*4}}{4\epsilon_{1}\epsilon_{3}(\tilde{\omega}-\epsilon_{1})(\tilde{\omega}-\epsilon_{3})}$ (58)
$\hskip-56.9055pt\times\Bigg[\frac{\tilde{C}_{0}(\alpha,\theta)}{|s_{0}|^{2}}+\frac{\tilde{C}_{2}(\alpha,\theta)+4\tilde{C}^{2}_{1}(\alpha,\theta)\cos^{2}\varphi}{|s_{\perp}|^{2}}$
$\hskip-56.9055pt-2{\rm Re}\frac{(2\epsilon_{1}-\tilde{\omega})(2\epsilon_{3}-\tilde{\omega})\tilde{C}_{1}(\alpha,\theta)\cos\varphi}{(s_{0}s^{*}_{\perp})}\Bigg],\qquad$
where using the shorthands $\cos\alpha=x_{\alpha}$ and $\cos\theta=x_{\theta}$,
$C_{0}$ $=$ $(2\epsilon_{1}^{2}-\epsilon_{1}\omega-p_{1}qx_{\alpha})(2\epsilon_{2}^{2}+\epsilon_{2}\omega+p_{2}qx_{\theta}),$ (59)
$C_{1}$ $=$ $p_{1}p_{2}(1-x_{\alpha}^{2})^{1/2}(1-x_{\theta}^{2})^{1/2},$ (60)
$C_{2}$ $=$ $2p_{1}^{2}(1-x_{\alpha}^{2})(\epsilon_{2}\omega-p_{2}qx_{\theta})$ (61)
$+$ $2p_{2}^{2}(1-x_{\theta}^{2})(p_{1}qx_{\alpha}-\epsilon_{1}\omega)$
$+$ $2(p_{1}qx_{\alpha}-\epsilon_{1}\omega)(\epsilon_{2}\omega-p_{2}qx_{\theta}),$
$\tilde{C}_{0}$ $=$ $(2\epsilon_{1}^{2}-\epsilon_{1}\tilde{\omega}-p_{1}\tilde{q}x_{\alpha})(2\epsilon_{3}^{2}-\epsilon_{3}\tilde{\omega}-p_{3}\tilde{q}x_{\theta}),$ (62)
$\tilde{C}_{1}$ $=$ $p_{1}p_{3}(1-x_{\alpha}^{2})^{1/2}(1-x_{\theta}^{2})^{1/2},$ (63)
[…]
$+$ $2(\epsilon_{1}\tilde{\omega}-p_{1}\tilde{q}x_{\alpha})(\epsilon_{3}\tilde{\omega}-p_{3}\tilde{q}x_{\theta}).$
III.2 The electron-ion and positron-ion scattering matrix elements
Figure 2: Diagrams describing the electron-ion (left) and positron-ion (right) scattering via exchange of a virtual plasmon.
In analogy to Eq. (50), the one-plasmon-exchange scattering amplitudes of electrons/positrons off ions can be written as (see Fig. 2)
${\cal M}_{12\to 34}^{ei}$ $=$ $-\frac{J_{0}\bar{J}_{0}}{t_{0}}+\frac{\bm{J}_{\perp}\cdot\bar{\bm{J}}_{\perp}}{t_{\perp}},$ (65)
${\cal M}_{12\to 34}^{pi}$ $=$ $-\frac{J^{\prime}_{0}\bar{J}_{0}}{t_{0}}+\frac{\bm{J}^{\prime}_{\perp}\cdot\bar{\bm{J}}_{\perp}}{t_{\perp}},$ (66)
where the electron, positron and ion 4-currents are given, respectively, by
$J^{\mu}=-e^{*}\bar{u}_{3}\gamma^{\mu}u_{1},\quad J^{\prime\mu}=-e^{*}\bar{v}_{1}\gamma^{\mu}v_{3},$ (67)
$\bar{J}^{\mu}=Z_{i}e^{*}(1,\bm{p}_{24}/M_{i}),$ (68)
and $\bm{J}_{\perp},\bm{J}^{\prime}_{\perp}$ are the components of the currents transverse to $\bm{q}$. Here $\bm{p}_{24}=(\bm{p}_{2}+\bm{p}_{4})/2$, therefore, as $\bm{p}_{4}=\bm{p}_{2}+\bm{q}$, we obtain $\bm{p}_{24}^{\perp}=\bm{p}_{2}^{\perp}=\bm{p}_{4}^{\perp}\equiv\bm{p}^{\prime\perp}$. Squaring these matrix elements and performing the spin-averaging procedure yields the final result (see Ref. Harutyunyan2016 )
$\frac{1}{2}\sum\limits_{\rm spins}|{\cal M}^{p,ei}_{12\to 34}|^{2}$ $=$ $\frac{Z_{i}^{2}e^{*4}}{2\epsilon_{1}(\epsilon_{1}-\omega)}$ (69)
$\hskip-71.13188pt\times\Biggl[\frac{\bar{C}_{0}(\alpha)}{|t_{0}|^{2}}-{\rm Re}\frac{2(2\epsilon_{1}-\omega)C_{1}(\alpha,\theta)\cos\varphi}{M_{i}(t_{0}t_{\perp}^{*})}$
$\hskip-28.45274pt+\frac{\bar{C}_{2}(\alpha,\theta)+2C^{2}_{1}(\alpha,\theta)\cos^{2}\varphi}{M^{2}_{i}|t_{\perp}|^{2}}\Biggr],$
with
$\bar{C}_{0}(\alpha)$ $=$ $2\epsilon_{1}^{2}-\epsilon_{1}\omega-p_{1}q\cos\alpha,$ (70)
$\bar{C}_{2}(\alpha,\theta)$ $=$ $p_{2}^{2}\sin^{2}\theta\,(p_{1}q\cos\alpha-\epsilon_{1}\omega).$ (71)
To account for ion–ion correlations and the finite size of the nuclei, the $ei$ and $pi$ matrix elements are multiplied by $S_{i}(q)F_{i}^{2}(q)$, where $S_{i}(q)$ is the ionic static structure factor and $F_{i}(q)$ is the nuclear form factor 1984ApJ…285..758I . (The effect of $F_{i}(q)$ on transport in the outer crust is, however, small.) Our numerical calculations are performed for a plasma consisting of a single ion species, using the fit formulas for the static structure factor $S(q)$ from Ref. Desbiens2016 .
For the polarization tensor, we use the hard-thermal-loop expressions of QED, see Ref. Harutyunyan2016 Eqs. (40) and (41). In the regime of low-frequency scattering $(x=\omega/q\ll 1)$, relevant for electron-ion and positron-ion collisions, the longitudinal and transverse components (assuming ultra-relativistic electrons and positrons) reduce to
$\Pi_{L}(q,\omega)$ $=q_{D}^{2}\chi_{l},$ $\Pi_{T}(q,\omega)$ $=q_{D}^{2}\chi_{t},$ (72)
with ${\rm Re}\chi_{l}=1-x^{2}$, ${\rm Im}\chi_{l}=-{\pi x}/{2}$, ${\rm Re}\chi_{t}=x^{2}$, and ${\rm Im}\chi_{t}={\pi x}/{4}.$ Here, $q_{D}$ is the Debye wave number, given by
$q_{D}^{2}=-\frac{4e^{2}}{\pi}\int_{0}^{\infty}d\epsilon\,p\,\epsilon\left(\frac{\partial f^{0+}}{\partial\epsilon}+\frac{\partial f^{0-}}{\partial\epsilon}\right).$ (73)
III.3 The collision rates
In this section, we summarize the results of computations of the collision rates given by Eqs. (39)–(42). The electron-ion collision rate was computed in Ref. Harutyunyan2016 with the assumption that the ion distribution is given in the classical Maxwell-Boltzmann form. This result can be generalized straightforwardly to obtain the electron-ion and positron-ion collision rates (39) and (40). We will use the simplified formula given by Eq. (32) of Ref. Harutyunyan2016 , which corresponds to the limit of static scattering. The final formulas read
$\nu_{pi/ei}$ $=$ $\frac{e^{4}Z_{i}^{2}n_{i}}{2\pi l^{\pm}}\!\int_{m}^{\infty}\!d\epsilon\,f^{0\pm}(\epsilon)\left[1-f^{0\pm}(\epsilon)\right]$ (74)
$\int_{0}^{2p}dq\,q^{3}S_{i}(q)F_{i}^{2}(q)\frac{4\epsilon^{2}-q^{2}}{|q^{2}+\Pi_{L}|^{2}}.$
The electron-positron collision rate is computed in Appendix B; the final result reads
$\nu_{ep}$ $=$ $\frac{e^{4}}{(2\pi)^{3}l^{-}}\!\int_{m}^{\infty}\!d\epsilon\!\int_{m}^{\infty}\!d\epsilon^{\prime}\!\int_{m-\epsilon^{\prime}}^{\epsilon-m}d\omega\,$ (75)
$\times$ $f^{0-}(\epsilon)f^{0+}(\epsilon^{\prime})\left[1-f^{0-}(\epsilon-\omega)\right]\left[1-f^{0+}(\epsilon+\omega)\right]$
$\times$ $\int_{Q_{-}}^{Q_{+}}dq\,\Bigg\{\frac{N_{L}}{4|q^{2}+\Pi_{L}|^{2}}+\frac{N_{T}}{2|q^{2}-\omega^{2}+\Pi_{T}|^{2}}\Bigg\}q^{2}$
$+$ $\frac{e^{4}}{(2\pi)^{3}l^{-}}\!\int_{m}^{\infty}\!d\epsilon\!\int_{m}^{\infty}\!d\epsilon^{\prime}\!\int_{\omega_{\rm min}}^{\infty}d{\omega}$
$\times$ $f^{0-}(\epsilon)f^{0+}({\omega}-\epsilon)\left[1-f^{0-}(\epsilon^{\prime})\right]\left[1-f^{0+}({\omega}-\epsilon^{\prime})\right]$
$\times$ $\int_{\tilde{Q}_{-}}^{\tilde{Q}_{+}}d{q}\,\Bigg\{\frac{u\tilde{N}_{L}}{4|{q}^{2}+{\Pi}_{L}|^{2}}+\frac{u\tilde{N}_{T}}{2|{q}^{2}-{\omega}^{2}+{\Pi}_{T}|^{2}}$
$+$ ${\rm Re}\frac{\tilde{N}_{LT}}{({q}^{2}+{\Pi}_{L})({q}^{2}-{\omega}^{2}+{\Pi}_{T})^{*}}\Bigg\},\qquad$
where the first and second terms represent the scattering and annihilation diagrams, respectively. Note that the variable $\epsilon$ stands for the energy of the initial state electron in both integrals, whereas the variable $\epsilon^{\prime}$ represents the energy of the final electron in the scattering integral and that of the initial positron in the annihilation integral. Thus, the variables $q$ and $\omega$ are the momentum and energy of the transferred plasmon in both cases, but the integration limits are different. The auxiliary functions in Eq. (75) are given by
$N_{L}$ $=$ $\big[(2\epsilon-\omega)^{2}-q^{2}\big]\big[(2\epsilon^{\prime}+\omega)^{2}-q^{2}\big],$ (76)
$N_{T}$ $=$ $\big[2p^{2}(1-x^{2})+q^{2}-\omega^{2}\big]\big[2p^{\prime 2}(1-y^{2})+q^{2}-\omega^{2}\big],$
$\tilde{N}_{L}$ $=$ $\big[(2\epsilon-\omega)^{2}-q^{2}\big]\big[(2\epsilon^{\prime}-\omega)^{2}-q^{2}\big],$ (78)
$\tilde{N}_{T}$ $=$ $\big[2p^{2}(1-{x}^{2})+{q}^{2}-{\omega}^{2}\big]\big[2p^{\prime 2}(1-z^{2})+{q}^{2}-{\omega}^{2}\big],$
$\tilde{N}_{LT}$ $=$ $2p^{2}p^{\prime 2}(2\epsilon-{\omega})(2\epsilon^{\prime}-{\omega})(1-{x}^{2})(1-z^{2}),$ (80)
where
$x$ $=$ $\frac{q^{2}-\omega^{2}+2\epsilon\omega}{2pq},\quad y=\frac{\omega^{2}-q^{2}+2\epsilon^{\prime}\omega}{2p^{\prime}q},$ (81)
$z$ $=$ $\frac{{q}^{2}-{\omega}^{2}+2\epsilon^{\prime}{\omega}}{2p^{\prime}{q}},\quad u=p^{2}+p^{\prime 2}-2pp^{\prime}xz.\quad$ (82)
The limits of the $q$-integration are given by $Q_{+}={\rm min}(q_{+},q^{\prime}_{+})$, $Q_{-}={\rm max}(q_{-},q^{\prime}_{-})$ with
$q_{\pm}$ $=$ $\left|\sqrt{(\epsilon-\omega)^{2}-m^{2}}\pm\sqrt{\epsilon^{2}-m^{2}}\right|,$ (83)
$q^{\prime}_{\pm}$ $=$ $\left|\sqrt{(\epsilon^{\prime}+\omega)^{2}-m^{2}}\pm\sqrt{\epsilon^{\prime 2}-m^{2}}\right|,$ (84)
[…]
$\tilde{q}_{\pm}={q}_{\pm},\quad\tilde{q}^{\prime}_{\pm}=\left|\sqrt{(\epsilon^{\prime}-{\omega})^{2}-m^{2}}\pm\sqrt{\epsilon^{\prime 2}-m^{2}}\right|.$ (85)
The lower bound for $\omega$ in the annihilation integral is $\omega_{\rm min}={\rm max}\{\epsilon,\epsilon^{\prime\prime}\}+m$, since the plasmon produced in the annihilation carries the total energy of the electron–positron pair.
Because the matrix element for $pe$ collisions is the same as the one for epep collisions, see Appendix A, we see from Eqs. (41) and (42) that the $pe$ collision rate can be obtained from the $ep$ collision rate by simply replacing $f^{0+}\leftrightarrow f^{0-}$ and $l^{-}\rightarrow l^{+}$, therefore
$\nu_{pe}$ $=$ $\frac{e^{4}}{(2\pi)^{3}l^{+}}\!\int_{m}^{\infty}\!d\epsilon\!\int_{m}^{\infty}\!d\epsilon^{\prime}\!\int_{m-\epsilon^{\prime}}^{\epsilon-m}d\omega\,f^{0+}(\epsilon)f^{0-}(\epsilon^{\prime})$ (86)
$\left[1-f^{0+}(\epsilon-\omega)\right]\left[1-f^{0-}(\epsilon+\omega)\right]$
$\times$ $\int_{Q_{-}}^{Q_{+}}dq\,\Bigg\{\frac{N_{L}}{4|q^{2}+\Pi_{L}|^{2}}+\frac{N_{T}}{2|q^{2}-\omega^{2}+\Pi_{T}|^{2}}\Bigg\}q^{2}$
$+$ $\frac{e^{4}}{(2\pi)^{3}l^{+}}\!\int_{m}^{\infty}\!d\epsilon\!\int_{m}^{\infty}\!d\epsilon^{\prime}\!\int_{\omega_{\rm min}}^{\infty}d{\omega}$
$f^{0+}(\epsilon)f^{0-}({\omega}-\epsilon)\left[1-f^{0+}(\epsilon^{\prime})\right]\left[1-f^{0-}({\omega}-\epsilon^{\prime})\right]$
$\times$ $\int_{\tilde{Q}_{-}}^{\tilde{Q}_{+}}d{q}\,\Bigg\{\frac{u\tilde{N}_{L}}{4|{q}^{2}+{\Pi}_{L}|^{2}}+\frac{u\tilde{N}_{T}}{2|{q}^{2}-{\omega}^{2}+{\Pi}_{T}|^{2}}$
$+$ ${\rm Re}\frac{\tilde{N}_{LT}}{({q}^{2}+{\Pi}_{L})({q}^{2}-{\omega}^{2}+{\Pi}_{T})^{*}}\Bigg\}.\qquad$
IV Numerical results
We performed numerical calculations of the electrical conductivities from Eq. (49) in the regime relevant to neutron star outer crusts, considering a plasma composed of a single ion species, $^{\makebox[7.97226pt][r]{$\scriptstyle 12$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{C}$}, across the entire temperature–density range studied. Below, we first present the phase structure of crustal matter, followed by a discussion of our new conductivity results. For practical astrophysical applications, all numerical values are given in c.g.s. units.
IV.1 Phase diagram of crustal plasma
Figure 3: The temperature-density phase diagram of crustal plasma composed of carbon. Ionic component of plasma forms a Boltzmann gas above the Coulomb temperature $T_{C}$, a classical liquid at $T_{p}\leq T\leq T_{C}$, and a quantum liquid below the plasma temperature $T_{p}$. Electrons become degenerate below the Fermi temperature $T_{F}$. The three curves around $T_{F}$ correspond to temperatures where the ratio of the total electron density $n^{-}$ to the net electron density $n_{e}$ reaches the value $r$.
To describe the state of the crustal plasma, we introduce the following physical quantities. The ion mass is defined as $M=Am_{n}$, where $A$ is the ion mass number, and $m_{n}\simeq 939$ MeV is the average nucleon mass. The ion number density $n_{i}$ is related to the density of matter $\rho$ by $n_{i}=\rho/M$, and we also define the radius of the spherical volume per ion (so-called the Wigner Seitz cell Shapiro1986 ) as $a=(4\pi n_{i}/3)^{-1/3}$.
The state of ions with charge number $Z$ is determined by the value of the Coulomb plasma parameter
$\Gamma=\frac{(Ze)^{2}}{aT},$ (87)
which is the ratio of the ion-ion Coulomb repulsion energy to their thermal energy. If $\Gamma\ll 1$ or, equivalently $T\gg T_{\rm C}\equiv(Ze)^{2}/a$, ions are weakly coupled and form a Boltzmann gas. In the regime where $\Gamma\geq 1$, ions are strongly coupled and form a liquid for values $\Gamma\leq\Gamma_{m}\simeq 160$ and a solid for $\Gamma>\Gamma_{m}$. The melting temperature of the lattice associated with $\Gamma_{m}$ is defined as $T_{m}=(Ze)^{2}/\Gamma_{m}a$. The plasma temperature
$T_{p}=\biggl(\frac{4\pi Z^{2}e^{2}n_{i}}{M}\biggl)^{1/2}$ (88)
determines the regime where the collective quantum effects in the plasma become important.
Figure 3 shows the temperature–density phase diagram of the crustal plasma composed of carbon $^{\makebox[7.97226pt][r]{$\scriptstyle 12$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{C}$}. The melting temperature lies below the plasma temperature and is therefore not shown in the diagram. This indicates that, as the temperature decreases, quantum effects in the carbon plasma become significant before the matter solidifies. The present study focuses on the portion of the phase diagram above the plasma temperature.
Electrons become degenerate below the Fermi temperature $T_{F}=\epsilon_{F}-m$, where the Fermi energy is defined as $\epsilon_{F}=(p_{F}^{2}+m^{2})^{1/2}$, and the Fermi momentum is given by $p_{F}=(3\pi^{2}n_{e})^{1/3}$. At finite temperatures, once the thermal energy of matter becomes sufficient to allow for creation of electron-positron pairs, a thermal population of positrons appears. The electron number excess over positrons $n_{e}$ is now the relevant quantity that determine the charge neutrality condition $n_{e}=Zn_{i}$. The electron chemical potential is found from the value implied by the charge conservation for electron excess $n_{e}=n^{-}-n^{+}$ for any given values of $\rho$ and $T$. The positron/electron proper number densities are given by
$n^{\pm}=\int\!\frac{2d\bm{p}}{(2\pi)^{3}}f^{0\pm}(\epsilon)=\frac{1}{\pi^{2}}\!\int_{m}^{\infty}\!d\epsilon\,p\epsilon f^{0\pm}(\epsilon),$ (89)
where the $f^{0\pm}$ are the Fermi distributions of electron and positrons introduced earlier. Clearly, $n^{-}>n_{e}$ when positrons’ population is not negligible.
In order to assess the relative importance of positrons in the plasma, we define the ratio $r=n^{-}/n_{e}$. Physically, $r$ is expected to rise rapidly with the temperature as a result of the fast opening of kinematic phase space for the pair creation with increasing $T$. This feature is clearly seen in Fig. 3, where we have plotted three lines which correspond to the values $r=2,5,10$. These curves highlight the crucial importance of including electron–positron pair creation processes in the composition of crustal matter, even at temperatures near the Fermi temperature. The ratio $r$ also increases with decreasing density at fixed $T$, as the pair creation rate is closely related to the ratio $T/T_{F}$. The shaded region at the top of the phase diagram corresponds to the regime where the number of created positrons, $n^{+}$, already exceeds the number of electrons in the cold $(T\ll T_{F})$ plasma. At temperatures $T\gg T_{F}$, the number of pairs increases approximately cubically with temperature, as implied by Eq. (89), yielding $n^{+}\simeq n^{-}\gg n_{e}$ in this regime.
IV.2 Electrical conductivity from electron-ion and positron-ion collisions
For clarity, we first consider the conductivities that include only electron–ion and positron–ion collisions. The full conductivities, which also account for electron–positron collisions, will be addressed in the next subsection. As indicated by Eqs. (43) and (74), the relaxation times for electrons and positrons are essentially identical in the regime where positrons are sufficiently abundant, since the integrals over combintations of Fermi functions in the numerators and denominators largely cancel. Consequently, the primary distinction between the electron and positron contributions to the electrical conductivity arises from the Fermi distribution functions in Eq. (49).
Let us first analyze the formula (49) in the limiting cases of strongly degenerate and nondegenerate electrons. In the degenerate limit where $T\ll T_{F}$, the positron contribution to the conductivity is suppressed, and $\sigma^{-}$ can be simplified via the substitution $\partial f^{0-}/\partial\epsilon\to-\delta(\epsilon-\epsilon_{F})$, which leads to $l^{\pm}\to p_{F}^{3}T/2\pi^{2}$ and, therefore (recall that $X^{\pm}=\tau^{\pm}/\epsilon$)
$\sigma\simeq\frac{n_{e}e^{2}\tau_{F}}{\epsilon_{F}},$ (90)
where $\tau_{F}\equiv\tau^{-}(\epsilon_{F})$. This is the well-known Drude formula for the electrical conductivity in the theory of metals Abrikosov:Fundamentals .
In the opposite limit of high temperatures $T\gg T_{F}$, we have also $T\gg\mu$, electrons and positrons are ultrarelativistic with the average thermal energy $\bar{\epsilon}\simeq 3T$, and their number densities are almost equal $n^{+}\simeq n^{-}\equiv\bar{n}\sim\bar{\epsilon}^{3}$. In this case $f^{0+}\simeq f^{0-}\equiv f^{0}\ll 1$, and Eq. (49) can be approximated as
$\sigma^{+}\simeq\sigma^{-}\simeq\frac{\bar{n}e^{2}\bar{\tau}}{\bar{\epsilon}},$ (91)
where $\bar{\tau}=\tau(\bar{\epsilon})$, and we used the definition of the statistical average of an energy-dependent quantity $F(\epsilon)$ as
$\bar{n}\bar{F}=\frac{1}{\pi^{2}}\int_{0}^{\infty}{p^{2}dp}\,F(\epsilon)f^{0}(\epsilon).$ (92)
Thus, the formulas for conductivity in both – strongly degenerate and nondegenerate regimes have the same form, but contain different energy scales, specifically $\epsilon_{F}$ in the degenerate regime and $\bar{\epsilon}\simeq 3T$ in the nondegenerate, ultrarelativistic regime.
The relaxation time (74) in the physical regimes of interest was discussed in detail in Ref. Harutyunyan2016 . On average, it was found that for carbon, $\tau$ varies with density and temperature according to the scaling $\tau(\epsilon)\propto\epsilon^{2}\rho^{-0.9}T^{-0.2}$. Using this scaling and the rough estimates $n_{e}\sim\epsilon_{F}^{3}$ and $\bar{n}\sim\bar{\epsilon}^{3}$, Eqs. (90) and (91) yield a universal scaling $\sigma\propto\epsilon^{4}\rho^{-0.9}T^{-0.2}$. This corresponds approximately to $\sigma\propto\rho^{0.5}T^{-0.2}$ in the degenerate regime and $\sigma\propto\rho^{-1}T^{4}$ in the nondegenerate regime.
Figure 4: Dependence of electronic (dashed lines) and positronic (dash-double-dotted lines) partial conductivities and their sum (solid lines) for three values of the temperature indicated in the plot. These are the conductivities resulting only from eiei and pipi collisions. The dotted lines show the conductivity from Ref. Harutyunyan2016 where positrons’ abundance in the matter is neglected.
Figure 5: The temperature dependence of electronic (dashed lines) and positronic (dash-double-dotted lines) partial conductivities and their sum (solid lines) for three values of the density indicated in the plot. These are the conductivities resulting only from eiei and pipi collisions. The dotted lines show the conductivity from Ref. Harutyunyan2016 where positrons’ abundance in the matter is neglected.
Figure 4 shows the partial conductivities $\sigma^{-}$ and $\sigma^{+}$, as well as the total conductivity $\sigma=\sigma^{-}+\sigma^{+}$ as functions of density for three temperature values, which are chosen to cover the range from the degenerate regime ($T=0.1$ MeV) to the nondegenerate regime ($T=10$ MeV). The intermediate value $T=1$ MeV is a representative of the transition between these regimes, which occurs around $\rho\simeq 10^{8}$ g cm-3, see Fig. 3. For comparison, we also plot the conductivities obtained previously in Ref. Harutyunyan2016 , where the presence of positrons in the matter was not taken into account.
The qualitative behavior of the conductivity in the degenerate regime is opposite to that in the transitional and nondegenerate regimes. This contrast can be understood by noting that at a temperature of $T=0.1$ MeV, positrons are effectively absent, since the temperature is far below the electron rest mass. The positron distribution is also strongly suppressed at $\rho\geq 10^{8}\,\mathrm{g\,cm^{-3}}$ for $T=1$ MeV due to Pauli blocking: as the Fermi energy increases, the number of kinematically accessible states for pair creation diminishes. In both situations, the present results for $\sigma$ coincide with the earlier work Harutyunyan2016 , as expected. In this regime, the conductivity scales as $\sigma\propto\rho^{0.5}$, because the increase in the electron density of states with $\rho$ outweighs the decrease in $\tau_{F}$.
Positrons are seen to make a significant contribution to the conductivity in the low-density regime $\rho\leq 10^{8}\,\mathrm{g\,cm^{-3}}$ for $T=1$ MeV, and across nearly the entire outer crust ($\rho\leq 10^{11}\,\mathrm{g\,cm^{-3}}$) for $T=10$ MeV. In this regime, the behavior of the conductivity reverses: at fixed temperature, it decreases with density, following the power law $\sigma\propto\tau\propto\rho^{-1}$. This occurs because the total number of electron–positron pairs, $\bar{n}$, is determined by temperature alone and is independent of the matter density.
The two conductivity branches representing the degenerate and nondegenerate regimes merge smoothly at a minimum located at their intersection. This transition point corresponds approximately to the temperature $T\simeq T_{F}/3$, where the two characteristic energy scales become comparable.
Let us now turn to the temperature dependence of the conductivity (see Fig. 5). The low-temperature branches again reproduce the results of Ref. Harutyunyan2016 , decreasing slowly with temperature due to the similar behavior of $\tau_{F}$, which is governed by ion–ion correlations. In Ref. Harutyunyan2016 , the conductivity was found to reach a minimum at the transition point $T\simeq T_{F}/3$, resulting from the reversed temperature dependence of the relaxation time in the nondegenerate regime.
As seen from Fig. 5, the position of the minimum around $T\simeq T_{F}/3$ remains unchanged also when positrons are present. However, the slope of the curves above the transition point rises abruptly, which is the consequence of the fast opening of the phase space for the pair creation with increasing temperature, as seen also from the phase diagram in Fig. 3. In contrast to the previous scaling $\sigma\propto T^{0.8}$ in the nondegenerate regime, we find here $\sigma\propto T^{4}$. To give a quantitative example, we compare the results with and without positrons at the lowest density considered here, $\rho=10^{6}\,\mathrm{g\,cm^{-3}}$. The inclusion of positrons increases the conductivity rougly by two orders of magnitude at $T=1$ MeV and by six orders of magnitude at $T=10$ MeV.
IV.3 Electrical conductivity including electron-positron collisions
Figure 6: Same as Fig. 4, but with inclusion of epep collisions.
We now discuss the electrical conductivities obtained in the full calculation, where electron–positron collisions are included as well. Let us recall that the electron–positron and positron–electron collision rates given by Eqs. (75) and (86) contain two distinct microscopic interaction channels: ordinary scattering and pair annihilation and creation processes. The latter is found to contribute significantly less to the total collision rate than the pure scattering channel.
Figure 6 shows the partial conductivities $\sigma^{-}$ and $\sigma^{+}$ and their sum $\sigma$ as functions of density for the same temperature values as in Fig. 4. Comparing these two figures, we see that the effect of electron-positron collisions becomes crucial in the low-density and high-temperature regime of non-degenerate matter where $T\gg T_{F}$, as expected. Here, the conductivity becomes density-independent. This behavior can be explained by the fact that $ep$ collisions heavily outweigh $ei$ and $pi$ collisions in this regime. As can be seen from Eq. (38), in this case the relaxation times contain only the collision rates $\nu_{ep}\simeq\nu_{pe}$, which are density independent in the regime $\mu_{e}\ll T$, thus resulting in density-independent conductivities in accordance with Eq. (91). In the degenerate regime, the electron-positron collisions are suppressed, and we recover the results of the previous section. Although the location of the minimum in the total conductivity remains unaffected by electron-positron collisions, these are non-negligible already at $T\geq T_{F}/3$.
Figure 7: Same as Fig. 5, but with inclusion of epep collisions.
Turning to the temperature dependence of conductivities shown in Fig. 7, we see again that the main difference as compared to Fig. 5 appears in the high temperature regime above the minimum point $T\simeq T_{F}/3$. At very high temperatures, where $T\gg T_{F}$, the conductivity $\sigma$ becomes essentially independent of density. On dimensional grounds, one then finds the scaling $\sigma\propto T$ in this regime, which agrees with the high-$T$ asymptotics shown in Fig. 7. This behavior is particularly evident at $\rho=10^{6}~\mathrm{g\,cm^{-3}}$. Immediately above the transition temperature $T\simeq T_{F}/3$, $\sigma$ follows the power-law dependence $\sigma\propto T^{4}$ as the collisions with ions are still dominant here. However, this fast increase is mitigated at higher temperatures as a result of intense electron-positron collisions, which surpass the $ei$ and $pi$ collision rates. As a numerical example, we find that at $\rho=10^{6}$ g cm-3 the ratios of conductivities we find to those computed earlier in Ref. Harutyunyan2016 are factor $20$ and $30$ for $T=1$ and $T=10$ MeV, respectively.
These factors decrease at higher densities but still reach nearly an order of magnitude across most of the outer crust at $T=10$ MeV. Concluding this section, we emphasize that the accounting for creation of positrons is essential for accurate determination of the transport properties in the outer crust of neutron stars, already at temperatures close to the Fermi temperature.
V Summary
In this work, we computed the electrical conductivity of hot QED plasma relevant to the outer crusts of neutron stars. The temperatures and densities studied here cover the transition from the degenerate to the nondegenerate regime for electrons. The novelty of our work is the inclusion of the contribution of positrons to the conductivity, which becomes increasingly important as the temperature increases. In this regime, the conductivity is dominated by thermally created electron-positron pairs which scatter off the correlated nuclei as well as off each other via screened electromagnetic force.
Solving a system of coupled Boltzmann kinetic equations for electrons and positrons, we expressed the conductivity in terms of electron and positron relaxation times. These contain electron-ion, positron-ion and electron-positron collisions rates, which were computed by using the relevant leading-order scattering diagrams which include all screening factors of strongly correlated crustal plasma.
Numerical results were obtained for conductivities for crustal matter composed of $^{\makebox[7.97226pt][r]{$\scriptstyle 12$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{C}$}. We found that the conductivity increases with density in the degenerate regime, in agreement with earlier results of Ref. Harutyunyan2016 , but in the nondegenerate regime it shows a decreasing behavior, in contrast to the results of previous studies. As a result, the conductivity crosses a minimum at the transition point from the nondegenerate to the degenerate regime. As a function of temperature, the conductivity grows with a power-law $\sigma\propto T^{4}$ in the semi-degenerate regime close to the transition temperature $T\simeq T_{F}/3$, because of intense creation of thermal electron-positron pairs. In this region, the number of pairs increases as $\propto T^{3}$, but the scattering is still mainly due to ions. At higher temperatures $T\gg T_{F}$, the electron-positron collisions become more important than electron-ion and positron-ion collisions, resulting in a slower, almost linear increase of $\sigma$ with temperature. We find that the inclusion of positrons enhances the electrical conductivity in the hot regions of the outer crust by factors of about 10. The results obtained in this work therefore, underscore the importance of accounting for positrons in the transport properties of heated plasma in neutron-star crusts.
Acknowledgements
The authors acknowledge support from the collaborative research grant No. 24RL-1C010 provided by the Higher Education and Science Committee (HESC) of the Republic of Armenia through the “Remote Laboratory” program. A. S. also acknowledges support from the Deutsche Forschungsgemeinschaft Grant No. SE 1836/6-1 and the Polish National Science Centre (NCN) Grant No. 2023/51/B/ST9/02798.
Appendix A The electron-positron scattering amplitude
In this Appendix, we will provide the details of the calculation of the spin-averaged electron-positron scattering amplitude. From Eqs. (50)–(III.1) we have
$\big|{\cal M}_{12\to 34}^{ep}\big|^{2}$ $=$ $\big|{\cal M}_{12\to 34}^{ep,t}\big|^{2}+\big|{\cal M}_{12\to 34}^{ep,s}\big|^{2}-2{\rm Re}\left({\cal M}_{12\to 34}^{ep,t}{\cal M}_{12\to 34}^{ep,s*}\right),$ (93)
$|{\cal M}^{ep,t}_{12\to 34}|^{2}$ $=$ $|{\cal M}^{t}_{L}|^{2}+|{\cal M}^{t}_{T}|^{2}-2{\rm Re}({\cal M}^{t}_{L}{\cal M}^{t*}_{T}),$ (94)
$|{\cal M}^{ep,s}_{12\to 34}|^{2}$ $=$ $|{\cal M}^{s}_{L}|^{2}+|{\cal M}^{s}_{T}|^{2}-2{\rm Re}({\cal M}^{s}_{L}{\cal M}^{s*}_{T}),$ (95)
${\cal M}_{12\to 34}^{ep,t}{\cal M}_{12\to 34}^{ep,s*}$ $=$ ${\cal M}^{t}_{L}{\cal M}^{s*}_{L}+{\cal M}^{t}_{T}{\cal M}^{s*}_{T}-({\cal M}^{t}_{L}{\cal M}^{s*}_{T}+{\cal M}^{t}_{T}{\cal M}^{s*}_{L}),$ (96)
with the $t$-channel (ordinary scattering) terms
$|{\cal M}^{t}_{L}|^{2}=\frac{J_{0}J_{0}^{*}J^{\prime}_{0}J_{0}^{\prime*}}{|t_{0}|^{2}},\quad|{\cal M}^{t}_{T}|^{2}=\frac{J_{i\perp}J_{k\perp}^{*}J_{i\perp}^{\prime}J_{k\perp}^{\prime*}}{|t_{\perp}|^{2}},\quad{\cal M}^{t}_{L}{\cal M}_{T}^{t*}=\frac{J_{0}J_{i\perp}^{*}J_{0}^{\prime}J_{i\perp}^{\prime*}}{t_{0}t_{\perp}^{*}},$ (97)
the $s$-channel (annihilation) terms
$|{\cal M}^{s}_{L}|^{2}=\frac{\tilde{J}_{0}\tilde{J}_{0}^{*}\tilde{J}_{0}^{\prime}\tilde{J}_{0}^{\prime*}}{|s_{0}|^{2}},\quad|{\cal M}^{s}_{T}|^{2}=\frac{\tilde{J}_{\perp i}\tilde{J}_{\perp k}^{*}\tilde{J}_{\perp i}^{\prime}\tilde{J}_{\perp k}^{\prime*}}{|s_{\perp}|^{2}},\quad{\cal M}^{s}_{L}{\cal M}_{T}^{s*}=\frac{\tilde{J}_{0}\tilde{J}_{\perp i}^{*}\tilde{J}_{0}^{\prime}\tilde{J}_{\perp i}^{\prime*}}{s_{0}s_{\perp}^{*}},$ (98)
and the interference terms that couple both channels
${\cal M}^{t}_{L}{\cal M}^{s*}_{L}=\frac{J_{0}\tilde{J}_{0}^{*}J^{\prime}_{0}\tilde{J}_{0}^{\prime*}}{t_{0}s_{0}^{*}},\qquad{\cal M}^{t}_{T}{\cal M}^{s*}_{T}=\frac{J_{i\perp}\tilde{J}_{k\perp}^{*}J_{i\perp}^{\prime}\tilde{J}_{k\perp}^{\prime*}}{t_{\perp}s_{\perp}^{*}},$ (99)
${\cal M}^{t}_{L}{\cal M}^{s*}_{T}=\frac{J_{0}\tilde{J}_{k\perp}^{*}J_{0}^{\prime}\tilde{J}_{k\perp}^{\prime*}}{t_{0}s_{\perp}^{*}},\qquad{\cal M}^{t}_{T}{\cal M}^{s*}_{L}=\frac{J_{i\perp}\tilde{J}_{0}^{*}J_{i\perp}^{\prime}\tilde{J}_{0}^{{}^{\prime}*}}{t_{\perp}s_{0}^{*}}.$ (100)
Equations (97) and (98) can be easily averaged over the spins, but the averaging of the interference terms (99) and (100) is quite cumbersome. Although all of these terms are generally of the same order of magnitude for relativistic particles, in this work, we will neglect, for simplicity, the interference terms and keep only the squares of the scattering and annihilation diagrams. Consider the $t$-channel first. The spin-averaged current-current couplings are given by
$\frac{1}{2}\sum\limits_{s_{1}s_{3}}J_{\mu}J_{\nu}^{*}$ $=$ $\frac{e^{*2}}{2\epsilon_{1}\epsilon_{3}}\Big[p_{1\mu}p_{3\nu}+p_{1\nu}p_{3\mu}-g_{\mu\nu}(\epsilon_{1}\epsilon_{3}-{\bm{p}}_{1}\cdot{\bm{p}}_{3}-m^{2})\Big].$ (101)
$\frac{1}{2}\sum\limits_{s_{2}s_{4}}J^{\prime}_{\mu}J^{\prime*}_{\nu}$ $=$ $\frac{e^{*2}}{2\epsilon_{2}\epsilon_{4}}\Big[p_{2\mu}p_{4\nu}+p_{2\nu}p_{4\mu}-g_{\mu\nu}(\epsilon_{2}\epsilon_{4}-{\bm{p}}_{2}\cdot{\bm{p}}_{4}-m^{2})\Big].$ (102)
Substituting the relevant component of the currents into Eq. (97) we obtain for the $t$-channel terms
$\frac{1}{4}\sum\limits_{\rm spins}|{\cal M}^{t}_{L}|^{2}$ $=$ $\frac{e^{*4}}{4\epsilon_{1}\epsilon_{2}\epsilon_{3}\epsilon_{4}}\frac{(\epsilon_{1}\epsilon_{3}+{\bm{p}}_{1}\cdot{\bm{p}}_{3}+m^{2})(\epsilon_{2}\epsilon_{4}+{\bm{p}}_{2}\cdot{\bm{p}}_{4}+m^{2})}{|t_{0}|^{2}},$ (103)
$\frac{1}{4}\sum\limits_{\rm spins}|{\cal M}^{t}_{T}|^{2}$ $=$ $\frac{e^{*4}}{4\epsilon_{1}\epsilon_{2}\epsilon_{3}\epsilon_{4}|t_{\perp}|^{2}}\Big[4({\bm{p}}^{\perp}\cdot{\bm{p}}^{\prime\perp})^{2}+2\left(\epsilon_{1}\epsilon_{3}-{\bm{p}}_{1}\cdot{\bm{p}}_{3}-m^{2}\right)\left(\epsilon_{2}\epsilon_{4}-{\bm{p}}_{2}\cdot{\bm{p}}_{4}-m^{2}\right)$ (104)
$+2({\bm{p}}^{\perp})^{2}\left(\epsilon_{2}\epsilon_{4}-{\bm{p}}_{2}\cdot{\bm{p}}_{4}-m^{2}\right)+2({\bm{p}}^{\prime\perp})^{2}\left(\epsilon_{1}\epsilon_{3}-{\bm{p}}_{1}\cdot{\bm{p}}_{3}-m^{2}\right)\Big],\quad$
$\frac{1}{4}\sum\limits_{\rm spins}{\cal M}^{t}_{L}{\cal M}_{T}^{t*}$ $=$ $\frac{e^{*4}}{4\epsilon_{1}\epsilon_{2}\epsilon_{3}\epsilon_{4}}\frac{({\bm{p}}^{\perp}\cdot{\bm{p}}^{\prime\perp})(\epsilon_{1}\epsilon_{2}+\epsilon_{1}\epsilon_{4}+\epsilon_{2}\epsilon_{3}+\epsilon_{3}\epsilon_{4})}{t_{0}t^{*}_{\perp}}.$ (105)
In Eqs. (104) and (105) we used the fact that $\delta^{\perp}_{ik}=\delta_{ik}-{q}_{i}{q}_{k}/{\bm{q}}^{2}$, and ${\bm{p}}_{3}^{\perp}={\bm{p}}_{1}^{\perp}\equiv{\bm{p}}^{\perp}$, ${\bm{p}}_{4}^{\perp}={\bm{p}}_{2}^{\perp}\equiv{\bm{p}}^{\prime\perp}$, which follow from the relations ${\bm{p}}_{3}={\bm{p}}_{1}-{\bm{q}}$, ${\bm{p}}_{4}={\bm{p}}_{2}+{\bm{q}}$.
We now modify the nominators using the substitutions ${\bm{p}}_{3}={\bm{p}}-{\bm{q}}$, ${\bm{p}}_{4}={\bm{p}}^{\prime}+{\bm{q}}$ and $\epsilon_{3}=\epsilon-\omega$, $\epsilon_{4}=\epsilon^{\prime}+\omega$ (with $p\equiv p_{1}$ and $p^{\prime}\equiv p_{2}$)
$\epsilon_{1}\epsilon_{3}+{\bm{p}}_{1}\cdot{\bm{p}}_{3}+m^{2}=2\epsilon^{2}-\epsilon\omega-{\bm{p}}\cdot{\bm{q}},$
$\epsilon_{2}\epsilon_{4}+{\bm{p}}_{2}\cdot{\bm{p}}_{4}+m^{2}=2\epsilon^{\prime 2}+\epsilon^{\prime}\omega+{\bm{p}}^{\prime}\cdot{\bm{q}},$
$\epsilon_{1}\epsilon_{3}-{\bm{p}}_{1}\cdot{\bm{p}}_{3}-m^{2}={\bm{p}}\cdot{\bm{q}}-\epsilon\omega,$
$\epsilon_{2}\epsilon_{4}-{\bm{p}}_{2}\cdot{\bm{p}}_{4}-m^{2}=\epsilon^{\prime}\omega-{\bm{p}}^{\prime}\cdot{\bm{q}},$
$({\bm{p}}^{\perp}\cdot{\bm{p}}^{\prime\perp})(\epsilon_{1}\epsilon_{2}+\epsilon_{1}\epsilon_{4}+\epsilon_{2}\epsilon_{3}+\epsilon_{3}\epsilon_{4})=({\bm{p}}^{\perp}\cdot{\bm{p}}^{\prime\perp})(2\epsilon-\omega)(2\epsilon^{\prime}+\omega).$
Substituting these expressions in Eqs. (103)–(105), for the spin-averaged square of the $t$-channel matrix element (94) we obtain
$\frac{1}{4}\sum\limits_{\rm spins}|{\cal M}^{ep,t}_{12\to 34}|^{2}$ $=$ $\frac{e^{*4}}{4\epsilon\epsilon^{\prime}(\epsilon-\omega)(\epsilon^{\prime}+\omega)}\Bigg\{\frac{(2\epsilon^{2}-\epsilon\omega-{\bm{p}}\cdot{\bm{q}})(2\epsilon^{\prime 2}+\epsilon^{\prime}\omega+{\bm{p}}^{\prime}\cdot{\bm{q}})}{|t_{0}|^{2}}$ (106)
$-2{\rm Re}\frac{({\bm{p}}^{\perp}\cdot{\bm{p}}^{\prime\perp})(2\epsilon-\omega)(2\epsilon^{\prime}+\omega)}{(t_{0}t^{*}_{\perp})}+\frac{1}{|t_{\perp}|^{2}}\Big[2\left({\bm{p}}\cdot{\bm{q}}-\epsilon\omega\right)\left(\epsilon^{\prime}\omega-{\bm{p}}^{\prime}\cdot{\bm{q}}\right)$
$+4({\bm{p}}^{\perp}\cdot{\bm{p}}^{\prime\perp})^{2}+2({\bm{p}}^{\perp})^{2}\left(\epsilon^{\prime}\omega-{\bm{p}}^{\prime}\cdot{\bm{q}}\right)+2({\bm{p}}^{\prime\perp})^{2}\left({\bm{p}}\cdot{\bm{q}}-\epsilon\omega\right)\Big]\Bigg\}.$
Figure 8: The relative orientation of three vectors $\bm{q}$}, $\bm{p}$} and $\bm{p}$$^{\prime}$ with angles defined by Eq. (107).
Defining the thee angles between the vectors $\bm{q}$}, $\bm{p}$} and $\bm{p}$$^{\prime}$ (see Fig. 8)
${\bm{p}}\cdot{\bm{q}}=pq\cos\alpha,\quad{\bm{p}}^{\prime}\cdot{\bm{q}}=p^{\prime}q\cos\theta,\quad{\bm{p}}_{\perp}\cdot{\bm{p}}^{\prime}_{\perp}=p_{\perp}p^{\prime}_{\perp}\cos\varphi,$ (107)
we obtain for Eq. (106)
$\frac{1}{4}\sum\limits_{\rm spins}|{\cal M}^{ep,t}_{12\to 34}|^{2}$ $=$ $\frac{e^{*4}}{4\epsilon\epsilon^{\prime}(\epsilon-\omega)(\epsilon^{\prime}+\omega)}\Bigg\{\frac{(2\epsilon^{2}-\epsilon\omega-pq\cos\alpha)(2\epsilon^{\prime 2}+\epsilon^{\prime}\omega+p^{\prime}q\cos\theta)}{|t_{0}|^{2}}$ (108)
$-2{\rm Re}\frac{(pp^{\prime}\sin\alpha\sin\theta\cos\varphi)(2\epsilon-\omega)(2\epsilon^{\prime}+\omega)}{(t_{0}t^{*}_{\perp})}$
$+\frac{1}{|t_{\perp}|^{2}}\Big[4(pp^{\prime}\sin\alpha\sin\theta\cos\varphi)^{2}+2\left(pq\cos\alpha-\epsilon\omega\right)\left(\epsilon^{\prime}\omega-p^{\prime}q\cos\theta\right)$
$+2p^{2}\sin^{2}\alpha\left(\epsilon^{\prime}\omega-p^{\prime}q\cos\theta\right)+2p^{\prime 2}\sin^{2}\theta\left(pq\cos\alpha-\epsilon\omega\right)\Big]\Bigg\}.$
The square of the annihilation diagram can be computed in an analogous manner. Performing the spin averaging of the relevant terms in Eq. (98), substituting Eq. (56), and using the same techniques as above, we obtain
$\frac{1}{2}\sum\limits_{s_{1}s_{2}}\tilde{J}_{\mu}\tilde{J}_{\nu}^{*}$ $=$ $\frac{e^{*2}}{2\epsilon_{1}\epsilon_{2}}\Big[p_{1\mu}p_{2\nu}+p_{1\nu}p_{2\mu}-g_{\mu\nu}(\epsilon_{1}\epsilon_{2}-{\bm{p}}_{1}\cdot{\bm{p}}_{2}+m^{2})\Big].$ (109)
$\frac{1}{2}\sum\limits_{s_{3}s_{4}}\tilde{J}^{\prime}_{\mu}\tilde{J}^{\prime*}_{\nu}$ $=$ $\frac{e^{*2}}{2\epsilon_{3}\epsilon_{4}}\Big[p_{3\mu}p_{4\nu}+p_{3\nu}p_{4\mu}-g_{\mu\nu}(\epsilon_{3}\epsilon_{4}-{\bm{p}}_{3}\cdot{\bm{p}}_{4}+m^{2})\Big],$ (110)
and
$\frac{1}{4}\sum\limits_{\rm spins}|{\cal M}^{ep,s}_{12\to 34}|^{2}$ $=$ $\frac{e^{*4}}{4\epsilon_{1}\epsilon_{2}\epsilon_{3}\epsilon_{4}}\Bigg\{\frac{(\epsilon_{1}\epsilon_{2}+{\bm{p}}_{1}\cdot{\bm{p}}_{2}-m^{2})(\epsilon_{3}\epsilon_{4}+{\bm{p}}_{3}\cdot{\bm{p}}_{4}-m^{2})}{|s_{0}|^{2}}$ (111)
$-2{\rm Re}\frac{({\bm{p}}_{1}^{\perp}\cdot{\bm{p}}_{3}^{\perp})(\epsilon_{1}\epsilon_{3}+\epsilon_{2}\epsilon_{4}-\epsilon_{1}\epsilon_{4}-\epsilon_{2}\epsilon_{3})}{(s_{0}s^{*}_{\perp})}$
$+\frac{1}{|s_{\perp}|^{2}}\Big[4({\bm{p}}_{1}^{\perp}\cdot{\bm{p}}_{3}^{\perp})^{2}+2\left(\epsilon_{1}\epsilon_{2}-{\bm{p}}_{1}\cdot{\bm{p}}_{2}+m^{2}\right)\left(\epsilon_{3}\epsilon_{4}-{\bm{p}}_{3}\cdot{\bm{p}}_{4}+m^{2}\right)$
$-2({\bm{p}}_{1}^{\perp})^{2}\left(\epsilon_{3}\epsilon_{4}-{\bm{p}}_{3}\cdot{\bm{p}}_{4}+m^{2}\right)-2({\bm{p}}_{3}^{\perp})^{2}\left(\epsilon_{1}\epsilon_{2}-{\bm{p}}_{1}\cdot{\bm{p}}_{2}+m^{2}\right)\Big]\Bigg\}.$
Defining $\epsilon^{\prime\prime}=\epsilon_{3}$, $p^{\prime\prime}=p_{3}$, $\tilde{\omega}=\epsilon+\epsilon_{2}=\epsilon^{\prime\prime}+\epsilon_{4}$, and $\tilde{{\bm{q}}}={\bm{p}}+{\bm{p}}_{2}={\bm{p}}^{\prime\prime}+{\bm{p}}_{4}$, and introducing three new angles by
${\bm{p}}\cdot\tilde{\bm{q}}=p\tilde{q}\cos\alpha^{\prime},\quad{\bm{p}}^{\prime\prime}\cdot\tilde{\bm{q}}=p^{\prime\prime}\tilde{q}\cos\theta^{\prime},\quad{\bm{p}}_{\perp}\cdot{\bm{p}}^{\prime\prime}_{\perp}=p_{\perp}p^{\prime\prime}_{\perp}\cos\varphi^{\prime},$ (112)
we obtain the final expression for the annihilation amplitude
$\frac{1}{4}\sum_{\rm spins}|{\cal M}^{ep,s}_{12\to 34}|^{2}$ $=$ $\frac{e^{*4}}{4\,\epsilon\,\epsilon^{\prime\prime}\,(\tilde{\omega}-\epsilon)(\tilde{\omega}-\epsilon^{\prime\prime})}\Bigg\{\frac{(2\epsilon^{2}-\epsilon\tilde{\omega}-p\tilde{q}\cos\alpha^{\prime})(2\epsilon^{\prime\prime 2}-\epsilon^{\prime\prime}\tilde{\omega}-p^{\prime\prime}\tilde{q}\cos\theta^{\prime})}{|s_{0}|^{2}}$ (113)
$-2\,\mathrm{Re}\frac{(p\,p^{\prime\prime}\,\sin\alpha^{\prime}\,\sin\theta^{\prime}\,\cos\varphi^{\prime})(2\epsilon-\tilde{\omega})(2\epsilon^{\prime\prime}-\tilde{\omega})}{s_{0}s_{\perp}^{*}}$
[…]
We next show that the matrix element for $pe$ collisions is the same as the one for $ep$ collisions. To check this, it is sufficient to interchange electrons and positrons, which corresponds to the following choice of Dirac 4-currents instead of Eqs. (55)–(56)
$J^{\mu}=-e^{*}\bar{v}_{1}\gamma^{\mu}v_{3},\quad J^{\prime\mu}=-e^{*}\bar{u}_{4}\gamma^{\mu}u_{2},$ (114)
$\tilde{J}^{\mu}=-e^{*}\bar{v}_{1}\gamma^{\mu}u_{2},\quad\tilde{J}^{\prime\mu}=-e^{*}\bar{u}_{4}\gamma^{\mu}v_{3}.$ (115)
Starting with these currents, a computation analogous to the one above leads then to exactly the same Eqs. (101) and (102), as well as Eqs. (109) and (110). Therefore, also the final forms of the scattering and annihilation terms in the matrix element are the same.
Appendix B Evaluating the electron-positron collision rate
Substituting the electron-positron scattering amplitude (93) in the collision rate (41) we obtain
$\nu_{ep}$ $=$ $\frac{(2\pi)^{-8}}{2l^{-}}\!\int\!d\bm{p}_{1}\,d\bm{p}_{2}\,d\bm{p}_{3}\,d\bm{p}_{4}\,\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4}){\bm{q}}^{2}$ (116)
$\times$ $f^{0-}_{1}f_{2}^{0+}(1-f^{0-}_{3})(1-f_{4}^{0+})\Big(|{\cal M}^{ep,t}_{12\to 34}|^{2}+|{\cal M}^{ep,s}_{12\to 34}|^{2}\Big),$
where we dropped the terms that mix the $t$ and $s$-channels. After introducing dummy integration over transferred momentum and energy for each channel separately, we write $\nu_{ep}=\nu^{t}_{ep}+\nu^{s}_{ep}$, where $\nu^{t}_{ep}$ and $\nu^{s}_{ep}$ denote the scattering and annihilation parts of the integrals, respectively; they are given by
$\nu^{t}_{ep}$ $=$ $\frac{(2\pi)^{-8}}{2l^{-}}\!\int\!d\omega\,d\bm{q}\!\int\!d\bm{p}\,d\bm{p}^{\prime}\,\delta(\epsilon-\epsilon_{3}-\omega)\delta(\epsilon^{\prime}-\epsilon_{4}+\omega)$ (117)
$\times{\bm{q}}^{2}f^{0-}_{1}f_{2}^{0+}(1-f^{0-}_{3})(1-f_{4}^{0+})|{\cal M}^{ep,t}_{12\to 34}|^{2},$
[…]
$\times{\bm{q}}^{2}f^{0-}_{1}f_{2}^{0+}(1-f^{0-}_{3})(1-f_{4}^{0+})|{\cal M}^{ep,s}_{12\to 34}|^{2}.$
Note that in Eq. (117), the energies $\epsilon_{3}$ and $\epsilon_{4}$ should be calculated for the momentum values ${\bm{p}}_{3}={\bm{p}}-{\bm{q}}$ and ${\bm{p}}_{4}={\bm{p}}^{\prime}+{\bm{q}}$, respectively, and in Eq. (118) $\epsilon_{2}$ and $\epsilon_{4}$ should be calculated for ${\bm{p}}_{2}=\tilde{{\bm{q}}}-{\bm{p}}$ and ${\bm{p}}_{4}=\tilde{{\bm{q}}}-{\bm{p}}^{\prime\prime}$. Here, we renamed $\bm{p}_{1}=\bm{p}$, $\bm{p}_{2}=\bm{p}^{\prime}$, $\bm{p}_{3}=\bm{p}^{\prime\prime}$, and $\epsilon_{1}=\epsilon$, $\epsilon_{2}=\epsilon^{\prime}$, $\epsilon_{3}=\epsilon^{\prime\prime}$. The four $\delta$-functions in Eqs. (117) and (118) can be written as
$\delta(\epsilon-\epsilon_{3}-\omega)$ $=$ $\delta\left(\epsilon-\omega-\sqrt{\epsilon^{2}+q^{2}-2pq\cos\alpha}\right)=\frac{\epsilon-\omega}{pq}\delta(\cos\alpha-x_{0})\theta(\epsilon-\omega),$ (119)
$\delta(\epsilon^{\prime}-\epsilon_{4}+\omega)$ $=$ $\delta\left(\epsilon^{\prime}+\omega-\sqrt{\epsilon^{\prime 2}+q^{2}+2p^{\prime}q\cos\theta}\right)=\frac{\epsilon^{\prime}+\omega}{p^{\prime}q}\delta(\cos\theta-y_{0})\theta(\epsilon^{\prime}+\omega),$ (120)
[…]
where we used the definitions of angles (107) and (112) and defined
$x_{0}=\frac{q^{2}-\omega^{2}+2\epsilon\omega}{2pq},\qquad y_{0}=\frac{\omega^{2}-q^{2}+2\epsilon^{\prime}\omega}{2p^{\prime}q},$ (123)
$\tilde{x}_{0}=\frac{\tilde{q}^{2}-\tilde{\omega}^{2}+2\epsilon\tilde{\omega}}{2p\tilde{q}},\qquad\tilde{y}_{0}=\frac{\tilde{q}^{2}-\tilde{\omega}^{2}+2\epsilon^{\prime\prime}\tilde{\omega}}{2p^{\prime\prime}\tilde{q}}$. (124)
Using these relations and substituting the matrix elements from Eqs. (57) and (58) into Eqs. (117) and (118), we obtain
$\nu^{t}_{ep}$ $=$ $\frac{(2\pi)^{-8}}{2l^{-}}4\pi\!\int_{m}^{\infty}\!\epsilon d\epsilon\!\int_{m}^{\infty}\!\epsilon^{\prime}d\epsilon^{\prime}\!\int_{-\epsilon^{\prime}}^{\epsilon}d\omega\,f^{0-}_{1}f_{2}^{0+}(1-f^{0-}_{3})(1-f_{4}^{0+})$ (125)
$\times 2\pi\!\int_{0}^{\infty}\!dq\!\int_{\pi}^{0}{d(\cos\alpha)}\delta(\cos\alpha-x_{0})I_{\Omega},$
$\nu^{s}_{ep}$ $=$ $\frac{(2\pi)^{-8}}{2l^{-}}4\pi\!\int_{m}^{\infty}\!\epsilon d\epsilon\!\int_{m}^{\infty}\!\epsilon^{\prime\prime}d\epsilon^{\prime\prime}\!\int_{\omega_{0}}^{\infty}d\tilde{\omega}\,f^{0-}_{1}f_{2}^{0+}(1-f^{0-}_{3})(1-f_{4}^{0+})$ (126)
$\times 2\pi\!\int_{0}^{\infty}\!d\tilde{q}\!\int_{\pi}^{0}{d(\cos\alpha^{\prime})}\delta(\cos\alpha^{\prime}-\tilde{x}_{0})\tilde{I}_{\Omega},$
where we used the relation $pdp=\epsilon d\epsilon$ and defined $\omega_{0}={\rm max}\{\epsilon,\epsilon^{\prime\prime}\}$. We also introduced the angular integrals as
$I_{\Omega}$ $=$ $\frac{e^{*4}}{4\epsilon\epsilon^{\prime}}\!\int_{\pi}^{0}d(\cos\theta)\delta(\cos\theta-y_{0})\!\int_{0}^{2\pi}\!d\varphi\Biggl[\frac{C_{0}(\alpha,\theta)}{|t_{0}|^{2}}+\frac{C_{2}(\alpha,\theta)+4C_{1}^{2}(\alpha,\theta)\cos^{2}\varphi}{|t_{\perp}|^{2}}$ (127)
$-2{\rm Re}\frac{(2\epsilon-\omega)(2\epsilon^{\prime}+\omega)C_{1}(\alpha,\theta)\cos\varphi}{t_{0}t^{*}_{\perp}}\Biggr]q^{2},$
$\tilde{I}_{\Omega}$ $=$ $\frac{e^{*4}}{4\epsilon\epsilon^{\prime\prime}}\!\int_{\pi}^{0}d(\cos\theta^{\prime})\delta(\cos\theta^{\prime}-\tilde{y}_{0})\!\int_{0}^{2\pi}\!d\varphi\Biggl[\frac{\tilde{C}_{0}(\alpha^{\prime},\theta^{\prime})}{|s_{0}|^{2}}+\frac{\tilde{C}_{2}(\alpha^{\prime},\theta^{\prime})+4\tilde{C}_{1}^{2}(\alpha^{\prime},\theta^{\prime})\cos^{2}\varphi^{\prime}}{|s_{\perp}|^{2}}$ (128)
$-2{\rm Re}\frac{(2\epsilon-\tilde{\omega})(2\epsilon^{\prime\prime}-\tilde{\omega})\tilde{C}_{1}(\alpha^{\prime},\theta^{\prime})\cos\varphi^{\prime}}{s_{0}s^{*}_{\perp}}\Biggr]q^{2}.$
We next carry out the azimuthal ($\varphi,\varphi^{\prime}$)-integrations in Eqs. (127) and (128) after substituting
${{\bm{q}}}^{2}=({\bm{p}}-{\bm{p}}^{\prime\prime})^{2}=p^{2}+p^{\prime\prime 2}-2pp^{\prime\prime}(\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos\varphi),$ (129)
where we used Eq. (107). Similar expression follows from Eq. (112) for the primed angles. The integrals of $\cos\varphi$ and $\cos^{3}\varphi$ over the interval $0,2\pi]$ vanish, while the integral of $\cos^{2}\varphi$ gives $\pi$ and the terms that do not depend on $\varphi$ integrate to $2\pi$. Therefore,
$I_{\Omega}$ $=$ $2\pi\frac{e^{*4}}{4\epsilon\epsilon^{\prime}}\!\int_{\pi}^{0}d(\cos\theta)\delta(\cos\theta-y_{0})\Biggl[\frac{C_{0}(\alpha,\theta)}{|t_{0}|^{2}}+\frac{C_{2}(\alpha,\theta)+2C_{1}^{2}(\alpha,\theta)}{|t_{\perp}|^{2}}\Biggr]q^{2},$ (130)
$\tilde{I}_{\Omega}$ $=$ $2\pi\frac{e^{*4}}{4\epsilon\epsilon^{\prime\prime}}\!\int_{\pi}^{0}d(\cos\theta)\delta(\cos\theta-\tilde{y}_{0})\Bigg\{\Biggl[\frac{\tilde{C}_{0}(\alpha,\theta)}{|s_{0}|^{2}}+\frac{\tilde{C}_{2}(\alpha,\theta)+2\tilde{C}_{1}^{2}(\alpha,\theta)}{|s_{\perp}|^{2}}\Biggr]$ (131)
$\times$ $\big(p^{2}+p^{\prime\prime 2}-2pp^{\prime\prime}\cos\alpha\cos\theta\big)+2{\rm Re}\frac{pp^{\prime\prime}(2\epsilon-\tilde{\omega})(2\epsilon^{\prime\prime}-\tilde{\omega})\tilde{C}_{1}(\alpha,\theta)\sin\alpha\sin\theta}{s_{0}s^{*}_{\perp}}\Bigg\},\qquad$
where we dropped primes on angles in the second expression, and
$\int_{\pi}^{0}{d(\cos\alpha)}\delta(\cos\alpha-x_{0})I_{\Omega}$ $=$ $2\pi\frac{e^{*4}}{4\epsilon\epsilon^{\prime}}\Biggl[\frac{{D}_{0}}{|t_{0}|^{2}}+\frac{{D}_{2}+2p^{2}p^{\prime 2}(1-x_{0}^{2})(1-y_{0}^{2})}{|t_{\perp}|^{2}}\Biggr]q^{2}\theta(1-|x_{0}|)\theta(1-|y_{0}|),$ (132)
$\int_{\pi}^{0}{d(\cos\alpha)}\delta(\cos\alpha-\tilde{x}_{0})\tilde{I}_{\Omega}$ $=$ $2\pi\frac{e^{*4}}{4\epsilon\epsilon^{\prime\prime}}\Bigg\{\Biggl[\frac{{\tilde{D}}_{0}}{|s_{0}|^{2}}+\frac{{\tilde{D}}_{2}+2p^{2}p^{\prime\prime 2}(1-\tilde{x}_{0}^{2})(1-\tilde{y}_{0}^{2})}{|s_{\perp}|^{2}}\Biggr]\big(p^{2}+p^{\prime\prime 2}-2pp^{\prime\prime}\tilde{x}_{0}\tilde{y}_{0}\big)$ (133)
$+$ $2{\rm Re}\frac{p^{2}p^{\prime\prime 2}(2\epsilon-\tilde{\omega})(2\epsilon^{\prime\prime}-\tilde{\omega})(1-\tilde{x}_{0}^{2})(1-\tilde{y}_{0}^{2})}{s_{0}s^{*}_{\perp}}\Bigg\}\theta(1-|\tilde{x}_{0}|)\theta(1-|\tilde{y}_{0}|).\qquad\qquad$
Here we used Eqs. (59)–(64) and defined new functions by
${D}_{0}$ $=$ $(2\epsilon^{2}-\epsilon\omega-pqx_{0})(2\epsilon^{\prime 2}+\epsilon^{\prime}\omega+p^{\prime}qy_{0}),$ (134)
${D}_{2}$ $=$ $2p^{2}(1-x_{0}^{2})\left(\epsilon^{\prime}\omega-p^{\prime}qy_{0}\right)+2p^{\prime 2}(1-y_{0}^{2})\left(pqx_{0}-\epsilon\omega\right)$ (135)
$+2\left(pqx_{0}-\epsilon\omega\right)\left(\epsilon^{\prime}\omega-p^{\prime}qy_{0}\right),$
${\tilde{D}}_{0}$ $=$ $(2\epsilon^{2}-\epsilon\tilde{\omega}-p\tilde{q}\tilde{x}_{0})(2\epsilon^{\prime\prime 2}-\epsilon^{\prime\prime}\tilde{\omega}-p^{\prime\prime}\tilde{q}\tilde{y}_{0}),$ (136)
[…]
$+2\left(\epsilon\tilde{\omega}-p\tilde{q}\tilde{x}_{0}\right)\left(\epsilon^{\prime\prime}\tilde{\omega}-p^{\prime\prime}\tilde{q}\tilde{y}_{0}\right).$
Substituting Eqs. (132) and (133) back into Eqs. (125) and (126), we obtain
$\nu^{t}_{ep}$ $=$ $(2\pi)^{-5}\frac{e^{*4}}{4l^{-}}\!\int_{m}^{\infty}\!d\epsilon\!\int_{m}^{\infty}\!d\epsilon^{\prime}\!\int_{-\epsilon^{\prime}}^{\epsilon}d\omega\,f^{0-}_{1}f_{2}^{0+}(1-f^{0-}_{3})(1-f_{4}^{0+})\int_{0}^{\infty}\!dq$ (138)
$\times\Biggl[\frac{{D}_{0}}{|t_{0}|^{2}}+\frac{{D}_{2}+2p^{2}p^{\prime 2}(1-x_{0}^{2})(1-y_{0}^{2})}{|t_{\perp}|^{2}}\Biggr]q^{2}\theta(1-|x_{0}|)\theta(1-|y_{0}|),$
$\nu^{s}_{ep}$ $=$ $(2\pi)^{-5}\frac{e^{*4}}{4l^{-}}\!\int_{m}^{\infty}\!d\epsilon\!\int_{m}^{\infty}\!d\epsilon^{\prime\prime}\!\int_{\omega_{0}}^{\infty}d\tilde{\omega}\,f^{0-}_{1}f_{2}^{0+}(1-f^{0-}_{3})(1-f_{4}^{0+})\int_{0}^{\infty}\!d\tilde{q}$ (139)
$\times\Bigg\{\Biggl[\frac{{\tilde{D}}_{0}}{|s_{0}|^{2}}+\frac{{\tilde{D}}_{2}+2p^{2}p^{\prime\prime 2}(1-\tilde{x}_{0}^{2})(1-\tilde{y}_{0}^{2})}{|s_{\perp}|^{2}}\Biggr]\big(p^{2}+p^{\prime\prime 2}-2pp^{\prime\prime}\tilde{x}_{0}\tilde{y}_{0}\big)$
$+2{\rm Re}\frac{p^{2}p^{\prime\prime 2}(2\epsilon-\tilde{\omega})(2\epsilon^{\prime\prime}-\tilde{\omega})(1-\tilde{x}_{0}^{2})(1-\tilde{y}_{0}^{2})}{s_{0}s^{*}_{\perp}}\Bigg\}\theta(1-|\tilde{x}_{0}|)\theta(1-|\tilde{y}_{0}|).$
The $\theta$-functions in Eq. (138) give two maxima and two minima for $q$: $_{-}\leq q\leq q_{+}, qq$$^{\prime}_{-}\leq q\leq q^{\prime}_{+}$, where
$q_{\pm}=\left|\sqrt{(\epsilon-\omega)^{2}-m^{2}}\pm\sqrt{\epsilon^{2}-m^{2}}\right|,\qquad q^{\prime}_{\pm}=\left|\sqrt{(\epsilon^{\prime}+\omega)^{2}-m^{2}}\pm\sqrt{\epsilon^{\prime 2}-m^{2}}\right|.$ (140)
Similarly, the $\theta$-functions in Eq. (139) imply $\tilde{q}_{-}\leq\tilde{q}\leq\tilde{q}_{+}$, $\tilde{q}^{\prime}_{-}\leq\tilde{q}\leq\tilde{q}^{\prime}_{+}$, with
$\tilde{q}_{\pm}=\left|\sqrt{(\epsilon-\tilde{\omega})^{2}-m^{2}}\pm\sqrt{\epsilon^{2}-m^{2}}\right|,\qquad\tilde{q}^{\prime}_{\pm}=\left|\sqrt{(\epsilon^{\prime\prime}-\tilde{\omega})^{2}-m^{2}}\pm\sqrt{\epsilon^{\prime\prime 2}-m^{2}}\right|.$ (141)
[…]
$\nu_{ep}^{t}$ $=$ $(2\pi)^{-5}\frac{e^{*4}}{4l^{-}}\!\int_{m}^{\infty}\!d\epsilon\!\int_{m}^{\infty}\!d\epsilon^{\prime}\!\int_{m-\epsilon^{\prime}}^{\epsilon-m}d\omega\,f^{0-}_{1}f_{2}^{0+}(1-f^{0-}_{3})(1-f_{4}^{0+})\int_{Q_{-}}^{Q_{+}}dq$
$\times$ $\Bigg[\frac{\big[(2\epsilon-\omega)^{2}-q^{2}\big]\big[(2\epsilon^{\prime}+\omega)^{2}-q^{2}\big]}{4|t_{0}|^{2}}+\frac{1}{|t_{\perp}|^{2}}\Big(p^{2}(1-x_{0}^{2})(q^{2}-\omega^{2})$
$+$ $p^{\prime 2}(1-y_{0}^{2})(q^{2}-\omega^{2})+\frac{1}{2}(q^{2}-\omega^{2})^{2}+2p^{2}p^{\prime 2}(1-x_{0}^{2})(1-y_{0}^{2})\Big)\Bigg]q^{2},$
$\nu_{ep}^{s}$ $=$ $(2\pi)^{-5}\frac{e^{*4}}{4l^{-}}\!\int_{m}^{\infty}\!d\epsilon\!\int_{m}^{\infty}\!d\epsilon^{\prime\prime}\!\int_{\omega_{0}+m}^{\infty}d\tilde{\omega}\,f^{0-}_{1}f_{2}^{0+}(1-f^{0-}_{3})(1-f_{4}^{0+})\int_{\tilde{Q}_{-}}^{\tilde{Q}_{+}}d\tilde{q}$ (142)
$\times$ $\Bigg\{\Bigg[\frac{\big[(2\epsilon-\tilde{\omega})^{2}-\tilde{q}^{2}\big]\big[(2\epsilon^{\prime\prime}-\tilde{\omega})^{2}-\tilde{q}^{2}\big]}{4|s_{0}|^{2}}+\frac{1}{|s_{\perp}|^{2}}\Big(p^{2}(1-\tilde{x}_{0}^{2})(\tilde{q}^{2}-\tilde{\omega}^{2})$
$+$ $p^{\prime\prime 2}(1-\tilde{y}_{0}^{2})(\tilde{q}^{2}-\tilde{\omega}^{2})+\frac{1}{2}(\tilde{q}^{2}-\tilde{\omega}^{2})^{2}+2p^{2}p^{\prime\prime 2}(1-\tilde{x}_{0}^{2})(1-\tilde{y}_{0}^{2})\Big)\Bigg]$
$\times$ $\big(p^{2}+p^{\prime\prime 2}-2pp^{\prime\prime}\tilde{x}_{0}\tilde{y}_{0}\big)+2{\rm Re}\frac{p^{2}p^{\prime\prime 2}(2\epsilon-\tilde{\omega})(2\epsilon^{\prime\prime}-\tilde{\omega})(1-\tilde{x}_{0}^{2})(1-\tilde{y}_{0}^{2})}{s_{0}s^{*}_{\perp}}\Bigg\},$
where we used Eqs. (123) and (124). Substituting $e^{*2}=4\pi e^{2}$, we obtain the final expression (75) for the electron-positron collision rate.

### v2

@@
 The relevant scattering matrix elements are calculated from one-plasmon exchange diagrams, with in-medium polarization tensors derived within the hard–thermal–loop effective theory.
-Numerical results are obtained for matter composed of carbon nuclei.
-We find that the conductivity rises with temperature, following a power law $\sigma\propto T^{4}$ in the semi-degenerate regime and $\sigma\propto T$ in the nondegenerate regime, due to the intense creation of thermal electron–positron pairs and the resulting collisions among them.
+Numerical results are obtained for matter composed of carbon, iron and helium nuclei.
+We find that the conductivity rises with temperature, following a power law $\sigma\propto T^{4}$ in the semidegenerate regime and $\sigma\propto T$ in the nondegenerate regime, due to the intense creation of thermal electron–positron pairs and the resulting collisions among them.
 These results highlight the importance of including positrons in the transport properties of heated, dense astrophysical plasmas.
 I Introduction
-Electrical conductivity of matter at subsaturation densities in neutron stars and cores of white dwarfs play a central role in several astrophysical problems, such as the magnetic field evolution 1992ApJ…395..250G ; 1995JApA…16..217B ; 1996ApJ…458..347M ; Ascenzi2024 , dissipation of magnetohydrodynamic waves 1979ApJ…227..995E ; 2014Ap…..57..530S , etc.
-Conductivity of dense neutron-star matter has been studied extensively in the cold regime relevant to isolated stars ($T\leq 1$ MeV) Flowers1976 ; Yakovlev1980 ; Urpin1980 ; Flowers1981 ; Itoh1983 ; Itoh1984 ; Nandkumar1984MNRAS ; Sedrakian1987 ; Itoh1993 ; Baiko1998 ; Potekhin1999 ; Shternin2006 ; Itoh2008 , see Ref.
+Electrical conductivity of matter at subsaturation densities in neutron stars and cores of white dwarfs plays a central role in several astrophysical problems, such as the magnetic field evolution 1992ApJ…395..250G ; 1995JApA…16..217B ; 1996ApJ…458..347M ; Ascenzi2024 , dissipation of magnetohydrodynamic waves 1979ApJ…227..995E ; 2014Ap…..57..530S , etc.
+The conductivity of dense neutron-star matter has been studied extensively in the cold regime relevant to isolated stars ($T\leq 1$ MeV) Flowers1976 ; Yakovlev1980 ; Urpin1980 ; Flowers1981 ; Itoh1983 ; Itoh1984 ; Nandkumar1984MNRAS ; Sedrakian1987 ; Itoh1993 ; Baiko1998 ; Potekhin1999 ; Shternin2006 ; Itoh2008 ; see Ref.
 Schmitt2018 for a recent review.
 However, formation of hot crustal matter in compact stars is anticipated by various astrophysical scenarios, such as supernova explosions leading to hot protoneutron stars, remnants of binary neutron star mergers, as well as heating of neutron star crusts in accreting binaries.
-In this regime, the plasma forms a liquid state of correlated ions, while the relativistic electron gas becomes non-degenerate.
+In this regime, the plasma forms a liquid state of correlated ions, while the relativistic electron gas becomes nondegenerate.
 This transition triggers intense production of electron–positron pairs, introducing a positron component into the plasma that contributes significantly to its conductivity.
-Finite temperature properties of such plasma are expected to significantly deviate from zero‑temperature (cold, degenerate) results.
+Finite temperature properties of such a plasma are expected to significantly deviate from zero‑temperature (cold, degenerate) results.
 The electrical conductivity of a one-component plasma in the hot regime under magnetic fields was investigated in Ref.
@@
 Harutyunyan2016 to a hot plasma state in which a non-negligible population of positrons coexists with electrons.
-Although positron contributions were neglected in earlier studies, they become critically important in the outer crust at temperatures $T\gtrsim 1\textrm{MeV}=1.16\times 10^{10}$ K.
+Although positron contributions were neglected in earlier studies, they become critically important in the outer crust of neutron stars at temperatures $T\gtrsim 1\textrm{MeV}=1.16\times 10^{10}$ K.
 In this regime, electrons and positrons constitute the primary charge carriers, and electrical conduction is governed mainly by electron–ion, positron–ion, and electron–positron scattering processes.
-To describe transport, we employ Boltzmann kinetic theory in the relaxation-time approximation to obtain the non-equilibrium distribution functions of electrons and positrons, fully accounting for electron–ion, positron–ion, and electron–positron collisions.
+To describe transport, we employ Boltzmann kinetic theory in the relaxation-time approximation to obtain the nonequilibrium distribution functions of electrons and positrons, fully accounting for electron–ion, positron–ion, and electron–positron collisions.
 In-medium screening effects are incorporated through the structure factor of the one-component classical plasma and the hard-thermal-loop effective-theory treatment of the plasmon self-energy Braaten1990 ; Braaten1992 ; Harutyunyan2016 .
@@
 II we derive the electrical conductivity for a multicomponent plasma consisting of electrons, positrons and ions, solving the system of linearized coupled Boltzmann equations for electron and positron distribution functions.
-In Section III, the matrix elements for electron-ion, positron-ion and positron-electron scattering processes and the resulting collision rates are discussed.
+In Sec.
+III, the matrix elements for electron-ion, positron-ion, and positron-electron scattering processes and the resulting collision rates are discussed.
 Section IV presents the phase structure of crustal matter and numerical results for electrical conductivity.
@@
 The kinetics of electrons and positrons is described by a system of coupled Boltzmann equations
-$\frac{\partial f^{-}}{\partial t}+\bm{v}\frac{\partial f^{-}}{\partial\bm{r}}-e(\bm{E}+\bm{v}\times\bm{H})\frac{\partial f^{-}}{\partial\bm{p}}=I^{-},$ (1)
-$\frac{\partial f^{+}}{\partial t}+\bm{v}\frac{\partial f^{+}}{\partial\bm{r}}+e(\bm{E}+\bm{v}\times\bm{H})\frac{\partial f^{+}}{\partial\bm{p}}=I^{+},$ (2)
-where $f^{-}$ and $f^{+}$ are the distribution functions for electrons and positrons, respectively, $\bm{E}$} and $\bm{H}$} are the electric and magnetic fields, $\bm{v}$} is the electron/positron velocity, and $I^{-}$ and $I^{+}$ are the collision integrals, respectively, for electrons and positrons, which include electron-ion and electron-positron scattering for $I^{-}$, and positron-ion and positron-electron scattering for $I^{+}$
+$\frac{\partial f^{-}}{\partial t}+\bm{v}\frac{\partial f^{-}}{\partial\bm{r}}-e(\bm{E}+\bm{v}\times\bm{H})\frac{\partial f^{-}}{\partial\bm{p}}=I^{-},\quad$ (1)
+$\frac{\partial f^{+}}{\partial t}+\bm{v}\frac{\partial f^{+}}{\partial\bm{r}}+e(\bm{E}+\bm{v}\times\bm{H})\frac{\partial f^{+}}{\partial\bm{p}}=I^{+},\quad$ (2)
+where $f^{-}$ and $f^{+}$ are the distribution functions for electrons and positrons, respectively, $\bm{E}$} and $\bm{H}$} are the electric and magnetic fields, $\bm{v}$} is the electron or positron velocity, and $I^{-}$ and $I^{+}$ are the collision integrals, respectively, for electrons and positrons, which include electron-ion and electron-positron scattering for $I^{-}$, and positron-ion and positron-electron scattering for $I^{+}$,
 $I^{-}=\sum_{i}I_{ei}+I_{ep},$ (3)
 $I^{+}=\sum_{i}I_{pi}+I_{pe},$ (4)
-where the sums in the first terms go over the type of ions (here we consider the general case of a multi-component plasma, where more than one sort of ions can be present simultaneously).
+where the sums in the first terms run over the ion species, allowing for the general case of a multicomponent plasma in which several types of ions may coexist.
+In general, the composition is described by a statistical distribution of nuclear species; however, in the present work, we assume that a single species is dominant, and we restrict our numerical computations to a plasma composed of this primary nuclear species.
+This approximation is valid in the low- to intermediate-temperature regime considered here, but may require corrections at higher temperatures, where differences in the binding energies of various nuclei become less significant in determining the composition of matter.
+A detailed study of a genuinely multicomponent plasma with temperature-dependent compositions is left for future work.
 For electron-ion and positron-ion collision integrals, we have
@@
 $\hskip-28.45274pt\times\Big[f^{+}_{1}g_{2}^{i}(1-f^{+}_{3})(1\pm g_{4}^{i})-f^{+}_{3}g_{4}^{i}(1-f^{+}_{1})(1\pm g_{2}^{i})\Big],\quad$
-where the indices $1$ and $3$ label the incoming and outgoing electrons/positrons, respectively, $g^{i}_{2,4}\equiv g^{i}(p_{2,4})$ are the ion distribution functions before and after the collision, and ${\cal M}_{12\to 34}^{ei,pi}$ is the electron/positron–ion scattering matrix element.
+where the indices 11 and $3$ label the incoming and outgoing electrons or positrons, respectively, $g^{i}_{2,4}\equiv g^{i}(p_{2,4})$ are the ion distribution functions before and after the collision, and ${\cal M}_{12\to 34}^{ei,pi}$ is the electron- or positron–ion scattering matrix element.
 The upper signs in Eqs.
@@
 $f^{\pm}=f^{0\pm}+\delta f^{\pm},\qquad\delta f^{\pm}=-\phi^{\pm}\frac{\partial f^{0\pm}}{\partial\epsilon},$ (9)
-where $f^{0\pm}({\epsilon})=[e^{\beta(\epsilon\pm\mu_{e})}+1]^{-1}$ are the local Fermi distribution functions for positrons/electrons with $\epsilon=\sqrt{p^{2}+m^{2}}$, $m$ is the electron mass, $\mu$ is the electron chemical potential, $\delta f^{\pm}\ll f^{0\pm}$ are small perturbations, and $\phi^{\pm}$ are unknown functions; additionally, introducing fermionic velocity $\bm{v}=\partial\epsilon/\partial\bm{p}=\bm{p}/\epsilon$, we quote some of the derivatives that will be used below
-$\frac{\partial f^{0\pm}}{\partial\bm{p}}$ $=$ $\bm{v}\frac{\partial f^{0\pm}}{\partial\epsilon},\quad\frac{\partial f^{0\pm}}{\partial\epsilon}=-\beta f^{0\pm}(1-f^{0\pm}).$ (10)
-We assume plasma is spatially uniform, time-independent and external fields are uniform and constant, i.e., there is no spatial variation of the fermionic distribution function, with only momentum space evolution.
+where $f^{0\pm}({\epsilon})=[e^{\beta(\epsilon\pm\mu_{e})}+1]^{-1}$ are the local Fermi distribution functions for positrons $(+)$ and electrons ()(-) with $\epsilon=\sqrt{p^{2}+m^{2}}$, $m$ is the electron mass, $\mu_{e}$ is the electron chemical potential, $\delta f^{\pm}\ll f^{0\pm}$ are small perturbations, and $\phi^{\pm}$ are unknown functions; additionally, introducing fermionic velocity $\bm{v}=\partial\epsilon/\partial\bm{p}=\bm{p}/\epsilon$, we give some of the derivatives that will be used below,
+$\frac{\partial f^{0\pm}}{\partial\bm{p}}$ $=$ $\bm{v}\frac{\partial f^{0\pm}}{\partial\epsilon},\quad\frac{\partial f^{0\pm}}{\partial\epsilon}=-\beta f^{0\pm}(1-f^{0\pm}).\quad$ (10)
+We assume plasma is spatially uniform, time independent and external fields are uniform and constant; i.e., there is no spatial variation of the fermionic distribution function, with only momentum space evolution.
 We next linearize the Boltzmann equation by substituting Eq.
@@
 (1) and (2) and keeping only the terms linear in $\delta f^{\pm}$ and in the electric field $\bm{E}$}, which implies that in the terms proportional to $\bm{E}$}, the substitutions $f^{\pm}\to f^{0\pm}$ can be made.
-We further neglect magnetic fields, which induce anisotropy in the conductivity; for the discussion of the effects of $\bm{H}$} field on transport, see Ref.
+We further neglect magnetic fields, which induce anisotropy in the conductivity; for the discussion of the effects of the $\bm{H}$} field on transport, see Ref.
 Harutyunyan2016 .
@@
 $\times$ $f_{1}^{0+}f_{2}^{0-}(1-f_{3}^{0+})(1-f_{4}^{0-}).\qquad$
-We seek solutions of Eqs. (11) and (12) in the form
+We search for the solutions of Eqs. (11) and (12) in the form
 $\phi^{\pm}={\pm}e\tau^{\pm}(\bm{v}\cdot\bm{E}),$ (17)
@@
 $I_{pe}$ $=$ $-e\bm{E}\sum_{pe}\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{-}_{2}\bm{v}_{2}-\tau^{+}_{3}\bm{v}_{3}+\tau^{-}_{4}\bm{v}_{4}\Big),\quad$ (21)
-and Eqs. (11) and (12) become
+and Eqs.
+(11) and (12) become (the cancellation of the vector $\bm{E}$ from both sides of the equations is justified as these should be satisfied for an arbitrary direction of $\bm{E}$; equivalently, one may project the equation along $\bm{E}$ and divide by $|\bm{E}|$)
 $-\frac{\partial f^{0-}}{\partial\epsilon}\bm{v}$ $=$ $\sum_{i}\sum_{ei}\Big(\tau^{-}_{1}\bm{v}_{1}-\tau^{-}_{3}\bm{v}_{3}\Big)$ (22)
-$\hskip-28.45274pt+\sum_{ep}\Big(\tau^{-}_{1}\bm{v}_{1}-\tau^{+}_{2}\bm{v}_{2}-\tau^{-}_{3}\bm{v}_{3}+\tau^{+}_{4}\bm{v}_{4}\Big).$
+$\hskip-28.45274pt+\sum_{ep}\Big(\tau^{-}_{1}\bm{v}_{1}-\tau^{+}_{2}\bm{v}_{2}-\tau^{-}_{3}\bm{v}_{3}+\tau^{+}_{4}\bm{v}_{4}\Big),\qquad$
 $-\frac{\partial f^{0+}}{\partial\epsilon}\bm{v}$ $=$ $\sum_{i}\sum_{pi}\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{+}_{3}\bm{v}_{3}\Big)$ (23)
-$\hskip-28.45274pt+\sum_{pe}\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{-}_{2}\bm{v}_{2}-\tau^{+}_{3}\bm{v}_{3}+\tau^{-}_{4}\bm{v}_{4}\Big).\quad$
+$\hskip-28.45274pt+\sum_{pe}\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{-}_{2}\bm{v}_{2}-\tau^{+}_{3}\bm{v}_{3}+\tau^{-}_{4}\bm{v}_{4}\Big).\qquad$
 We multiply Eqs. (22) and (23) by $\tau_{1}^{-}{\bm{v}}_{1}$ and $\tau_{1}^{+}{\bm{v}}_{1}$, respectively, and integrate over ${\bm{p}}_{1}\equiv{\bm{p}}$, yielding,
@@
 $+$ $\frac{1}{2}\sum_{1}\sum_{ep}\Big(\tau_{1}^{-}\bm{v}_{1}-\tau^{-}_{3}\bm{v}_{3}\Big)$
-$\hskip-31.2982pt\Big(\tau^{-}_{1}\bm{v}_{1}-\tau^{+}_{2}\bm{v}_{2}-\tau^{-}_{3}\bm{v}_{3}+\tau^{+}_{4}\bm{v}_{4}\Big),$
+$\hskip-31.2982pt\times\Big(\tau^{-}_{1}\bm{v}_{1}-\tau^{+}_{2}\bm{v}_{2}-\tau^{-}_{3}\bm{v}_{3}+\tau^{+}_{4}\bm{v}_{4}\Big),\qquad$
 $-\sum_{1}\frac{\partial f^{0+}}{\partial\epsilon}\tau^{+}{\bm{v}}^{2}$ $=$ $\frac{1}{2}\sum_{1,i}\sum_{pi}\Big(\tau_{1}^{+}\bm{v}_{1}-\tau_{3}^{+}\bm{v}_{3}\Big)^{2}$ (25)
 $+$ $\frac{1}{2}\sum_{1}\sum_{pe}\Big(\tau_{1}^{+}\bm{v}_{1}-\tau^{+}_{3}\bm{v}_{3}\Big)$
-$\hskip-34.14322pt\times\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{-}_{2}\bm{v}_{2}-\tau^{+}_{3}\bm{v}_{3}+\tau^{-}_{4}\bm{v}_{4}\Big),\quad$
+$\hskip-34.14322pt\times\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{-}_{2}\bm{v}_{2}-\tau^{+}_{3}\bm{v}_{3}+\tau^{-}_{4}\bm{v}_{4}\Big),\qquad$
 where we performed permutations $1\leftrightarrow 3$ and $2\leftrightarrow 4$.
@@
 $\nu_{ep}$ $=$ $\bigg[-\sum_{1}\frac{\partial f^{0-}}{\partial\epsilon}\frac{p^{2}}{\epsilon}\bigg]^{-1}\frac{1}{2}\sum_{1}\sum_{ep}{\bm{q}}^{2}.$ (32)
-Similar expressions hold for positons with the replacements $ei\to pi$, $ep\to pe$ and $f^{0-}\to f^{0+}$.
+Similar expressions hold for positrons with the replacements $ei\to pi$, $ep\to pe$ and $f^{0-}\to f^{0+}$.
 These quantities represent appropriately weighted mean collision rates between the corresponding particles.
-Solving (29) and (30) then yields
+Solving Eqs.
+(29) and (30) then yields
 $X^{-}$ $=$ $\frac{\sum_{i}\nu_{pi}+\nu_{pe}-\nu_{ep}}{\big(\sum_{i}\nu_{ei}+\nu_{ep}\big)\big(\sum_{i}\nu_{pi}+\nu_{pe}\big)-\nu_{ep}\nu_{pe}},$ (33)
@@
 $(X^{-})^{-1}=\sum_{i}\nu_{ei},\qquad(X^{+})^{-1}=\sum_{i}\nu_{pi},$ (35)
-where the collision rates on different types of nuclei sum up.
-In the regimes where electron-positron collisions become important, the number of electrons and positrons is almost equal, and we find that practically always $\nu_{ep}\simeq\nu_{pe}$, such that $\nu_{ep}-\nu_{pe}\ll\nu_{ei}\simeq\nu_{pi}$.
-Then Eqs.
+where the collision rates on different types of nuclei are summed up.
+In the regimes where electron-positron collisions become important, the numbers of electrons and positrons are almost equal, and we find that, practically always, $\nu_{ep}\simeq\nu_{pe}$, such that $\nu_{ep}-\nu_{pe}\ll\nu_{ei}\simeq\nu_{pi}$.
+Then, Eqs.
 (33) and (34) can be simplified to
-$X^{-}$ $=$ $\frac{\sum_{i}\nu_{pi}}{\big(\sum_{i}\nu_{ei}\big)\big(\sum_{i}\nu_{pi}\big)+\nu_{ep}\sum_{i}(\nu_{ei}+\nu_{pi})},$ (36)
-$X^{+}$ $=$ $\frac{\sum_{i}\nu_{ei}}{\big(\sum_{i}\nu_{ei}\big)\big(\sum_{i}\nu_{pi}\big)+\nu_{ep}\sum_{i}(\nu_{ei}+\nu_{pi})}.$ (37)
-In the opposite limit of fast electron-positron scattering we have $\nu_{ei}\simeq\nu_{pi}\ll\nu_{ep}\simeq\nu_{pe}$, therefore
+$X^{-}$ $=$ $\frac{\sum_{i}\nu_{pi}}{\big(\sum_{i}\nu_{ei}\big)\big(\sum_{i}\nu_{pi}\big)+\nu_{ep}\sum_{i}(\nu_{ei}+\nu_{pi})},\quad$ (36)
+$X^{+}$ $=$ $\frac{\sum_{i}\nu_{ei}}{\big(\sum_{i}\nu_{ei}\big)\big(\sum_{i}\nu_{pi}\big)+\nu_{ep}\sum_{i}(\nu_{ei}+\nu_{pi})}.\quad$ (37)
+In the limit of fast electron-positron scattering we have also $\nu_{ei}\simeq\nu_{pi}\ll\nu_{ep}\simeq\nu_{pe}$; therefore,
 $X^{-}=X^{+}=\frac{1}{2\nu_{ep}}=\frac{1}{2\nu_{pe}},$ (38)
-which can be derived also directly from Eqs. (29) and (30).
-We next write explicit expressions for the mean collision rates by substituting Eqs. (13)–(16) into (31) and (32) (and their analogues for positrons) to obtain
+which can also be derived directly from Eqs. (29) and (30).
+We next write explicit expressions for the mean collision rates by substituting Eqs. (13)–(16) into (31) and (32) (and their analogs for positrons) to obtain
 $\nu_{ei}$ $=$ $\frac{(2\pi)^{4}}{2l^{-}}\!\sum\limits_{1234}|{\cal M}_{12\to 34}^{ei}|^{2}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})$ (39)
@@
 $j^{+}_{k}$ $=$ $\int\!\frac{2d\bm{p}}{(2\pi)^{3}}ev_{k}\,\delta f^{+}=\sigma^{+}_{kj}E_{j},$ (45)
-where we used Eqs. (9) and (17), and defined the conductivity matrices (the factor 2 accounts for spin degeneracy) as
+where we used Eqs. (9) and (17) and defined the conductivity matrices (the factor 2 accounts for spin degeneracy) as
 $\sigma^{\pm}_{kj}=-\int\!\frac{2d\bm{p}}{(2\pi)^{3}}\frac{\partial f^{0\pm}}{\partial\epsilon}e^{2}\tau^{\pm}v_{k}v_{j}=\delta_{kj}\sigma^{\pm},$ (46)
@@
 The electrical conductivity is thus fully determined once the relaxation times $\tau^{\pm}$ are known.
-Employing the ansatz $\tau^{\pm}=X^{\pm}\epsilon$, Eq.
+By employing the ansatz $\tau^{\pm}=\epsilon X^{\pm}$, Eq.
 (47) becomes
@@
 The plasmon self-energy is given by the polarization tensor $\Pi_{\mu\nu}(\omega,{\bm{q}})$ shown by the closed loops.
+We remark that the range of validity of the Boltzmann kinetic theory employed here can be derived from the common considerations Lifshitz_Pitaevskii1981 .
+The main assumption is that the collisions are instantaneous, which is justified if the average scattering length which has the order of the Debye screening radius $r_{D}=q_{D}^{-1}$ of the electric field (see Sec.
+III.2 for details) is much smaller than the electron or positron mean free paths $l^{\pm}\simeq\tau^{\pm}$ between collisions (assuming ultra-relativistic particles with average velocities $\bar{v}^{\pm}\simeq 1$).
+The particle mean collision times and related mean free paths are supposed to be much smaller than the characteristic length scale $L$ and the inverse frequency $\Omega^{-1}$ of macroscopic perturbations of the electric field.
+Thus, the general validity condition can be written as
+$q_{D}^{-1}\ll\tau^{\pm}\ll{\rm min}\{L,\Omega^{-1}\}.$
+Our numerical results (see Sec.
+IV, Figs.
+46) show that in the regime of interest $0.03\leq q_{D}\leq 2$ MeV and $10^{2}\leq\tau^{\pm}\leq 10^{4}$ MeV-1 (recall that 1 s= $1.52\times 10^{21}$ MeV-1); therefore the condition $q_{D}\tau^{\pm}\gg 1$ is satisfied practically in the whole regime of interest with high accuracy.
 III Scattering amplitudes and collision rates
@@
 In addition, we will also discuss the electron-positron collision rates using the obtained expressions for the matrix elements.
-III.1 The electron-positron scattering matrix element
+III.1 Electron-positron scattering matrix element
 The screened electron–positron scattering matrix element can be expressed as the sum of two contributions, corresponding to proper scattering and annihilation processes (analogous to Bhabha scattering in vacuum); see Fig.
@@
 $J^{\mu}=-e^{*}\bar{u}_{3}\gamma^{\mu}u_{1},\quad J^{\prime\mu}=-e^{*}\bar{v}_{2}\gamma^{\mu}v_{4},$ (55)
-$\tilde{J}^{\mu}=-e^{*}\bar{v}_{2}\gamma^{\mu}u_{1},\quad\tilde{J}^{\prime\mu}=-e^{*}\bar{u}_{3}\gamma^{\mu}v_{4},$ (56)
-where $e^{*}=\sqrt{4\pi}e$, $(\omega,{\bm{q}})=p_{1}-p_{3}=p_{4}-p_{2}$, $(\tilde{\omega},\tilde{{\bm{q}}})=p_{1}+p_{2}=p_{3}+p_{4}$, $u_{a}={u}^{s_{a}}(p_{a})$, $v_{b}={v}^{s_{b}}(p_{b})$, $\bm{J}_{\perp},\bm{J}^{\prime}_{\perp}$ are the components of these currents transverse to $\bm{q}$, and $\tilde{\bm{J}}_{\perp},\tilde{\bm{J}}^{\prime}_{\perp}$ are those transverse to $\tilde{\bm{q}}$.
+$\tilde{J}^{\mu}=-e^{*}\bar{v}_{2}\gamma^{\mu}u_{1},\quad\tilde{J}^{\prime\mu}=-e^{*}\bar{u}_{3}\gamma^{\mu}v_{4}.$ (56)
+Here, $e^{*}=\sqrt{4\pi}e$, $(\omega,{\bm{q}})=p_{1}-p_{3}=p_{4}-p_{2}$, $(\tilde{\omega},\tilde{{\bm{q}}})=p_{1}+p_{2}=p_{3}+p_{4}$, $u_{a}={u}^{s_{a}}(p_{a})$, $v_{b}={v}^{s_{b}}(p_{b})$, $\bm{J}_{\perp},\bm{J}^{\prime}_{\perp}$ are the components of these currents transverse to $\bm{q}$, and $\tilde{\bm{J}}_{\perp},\tilde{\bm{J}}^{\prime}_{\perp}$ are those transverse to $\tilde{\bm{q}}$.
 The screening of the interaction is taken into account in terms of the longitudinal $\Pi_{L}(\omega,q)$ and transverse $\Pi_{T}(\omega,q)$ components of the polarization tensor.
-Details of the squaring and spin-averaging of the $ep$ matrix element are given in Appendix A.
-In this work, we neglect the interference terms between the $t$- and $s$-channels in the electron–positron scattering amplitude.
+Figure 2: Diagrams describing the electron-ion (left) and positron-ion (right) scattering via exchange of a virtual plasmon.
+Details of the squaring and spin averaging of the $ep$ matrix element are given in Appendix A.
+In this work, for simplicity, we neglect the interference terms between the $t$- and $s$-channels in the electron–positron scattering amplitude.
 The final results for the $t$- and $s$-channel spin-averaged scattering amplitudes are
@@
 $\hskip-56.9055pt-2{\rm Re}\frac{(2\epsilon_{1}-\tilde{\omega})(2\epsilon_{3}-\tilde{\omega})\tilde{C}_{1}(\alpha,\theta)\cos\varphi}{(s_{0}s^{*}_{\perp})}\Bigg],\qquad$
-where using the shorthands $\cos\alpha=x_{\alpha}$ and $\cos\theta=x_{\theta}$,
+where, using the shorthands $\cos\alpha=x_{\alpha}$ and $\cos\theta=x_{\theta}$,
 $C_{0}$ $=$ $(2\epsilon_{1}^{2}-\epsilon_{1}\omega-p_{1}qx_{\alpha})(2\epsilon_{2}^{2}+\epsilon_{2}\omega+p_{2}qx_{\theta}),$ (59)
@@
 $+$ $2(\epsilon_{1}\tilde{\omega}-p_{1}\tilde{q}x_{\alpha})(\epsilon_{3}\tilde{\omega}-p_{3}\tilde{q}x_{\theta}).$
-III.2 The electron-ion and positron-ion scattering matrix elements
-Figure 2: Diagrams describing the electron-ion (left) and positron-ion (right) scattering via exchange of a virtual plasmon.
-In analogy to Eq. (50), the one-plasmon-exchange scattering amplitudes of electrons/positrons off ions can be written as (see Fig. 2)
+III.2 Electron-ion and positron-ion scattering matrix elements
+In analogy to Eq. (50), the one-plasmon-exchange scattering amplitudes of electrons or positrons off ions can be written as (see Fig. 2)
 ${\cal M}_{12\to 34}^{ei}$ $=$ $-\frac{J_{0}\bar{J}_{0}}{t_{0}}+\frac{\bm{J}_{\perp}\cdot\bar{\bm{J}}_{\perp}}{t_{\perp}},$ (65)
 ${\cal M}_{12\to 34}^{pi}$ $=$ $-\frac{J^{\prime}_{0}\bar{J}_{0}}{t_{0}}+\frac{\bm{J}^{\prime}_{\perp}\cdot\bar{\bm{J}}_{\perp}}{t_{\perp}},$ (66)
-where the electron, positron and ion 4-currents are given, respectively, by
+where the electron, positron, and ion 4-currents are given, respectively, by
 $J^{\mu}=-e^{*}\bar{u}_{3}\gamma^{\mu}u_{1},\quad J^{\prime\mu}=-e^{*}\bar{v}_{1}\gamma^{\mu}v_{3},$ (67)
@@
 and $\bm{J}_{\perp},\bm{J}^{\prime}_{\perp}$ are the components of the currents transverse to $\bm{q}$.
-Here $\bm{p}_{24}=(\bm{p}_{2}+\bm{p}_{4})/2$, therefore, as $\bm{p}_{4}=\bm{p}_{2}+\bm{q}$, we obtain $\bm{p}_{24}^{\perp}=\bm{p}_{2}^{\perp}=\bm{p}_{4}^{\perp}\equiv\bm{p}^{\prime\perp}$.
+Here, $\bm{p}_{24}=(\bm{p}_{2}+\bm{p}_{4})/2$; therefore, as $\bm{p}_{4}=\bm{p}_{2}+\bm{q}$, we obtain $\bm{p}_{24}^{\perp}=\bm{p}_{2}^{\perp}=\bm{p}_{4}^{\perp}\equiv\bm{p}^{\prime\perp}$.
 Squaring these matrix elements and performing the spin-averaging procedure yields the final result (see Ref.
-Harutyunyan2016 )
+Harutyunyan2016 ),
 $\frac{1}{2}\sum\limits_{\rm spins}|{\cal M}^{p,ei}_{12\to 34}|^{2}$ $=$ $\frac{Z_{i}^{2}e^{*4}}{2\epsilon_{1}(\epsilon_{1}-\omega)}$ (69)
 $\hskip-71.13188pt\times\Biggl[\frac{\bar{C}_{0}(\alpha)}{|t_{0}|^{2}}-{\rm Re}\frac{2(2\epsilon_{1}-\omega)C_{1}(\alpha,\theta)\cos\varphi}{M_{i}(t_{0}t_{\perp}^{*})}$
-$\hskip-28.45274pt+\frac{\bar{C}_{2}(\alpha,\theta)+2C^{2}_{1}(\alpha,\theta)\cos^{2}\varphi}{M^{2}_{i}|t_{\perp}|^{2}}\Biggr],$
+$\hskip-28.45274pt+\frac{\bar{C}_{2}(\alpha,\theta)+2C^{2}_{1}(\alpha,\theta)\cos^{2}\varphi}{M^{2}_{i}|t_{\perp}|^{2}}\Biggr],\quad$
 with
@@
 Desbiens2016 .
-For the polarization tensor, we use the hard-thermal-loop expressions of QED, see Ref.
-Harutyunyan2016 Eqs.
+For the polarization tensor, we use the hard-thermal-loop expressions of QED; see Ref.
+Harutyunyan2016 , Eqs.
 (43) and (44).
-In the regime of low-frequency scattering $(x=\omega/q\ll 1)$, relevant for electron-ion and positron-ion collisions, the longitudinal and transverse components (assuming ultra-relativistic electrons and positrons) reduce to
-$\Pi_{L}(q,\omega)$ $=q_{D}^{2}\chi_{l},$ $\Pi_{T}(q,\omega)$ $=q_{D}^{2}\chi_{t},$ (72)
+This approximation assumes that $\omega\ll\epsilon$ and $q\ll p$; i.e., the 4-momentum of the plasmon is much smaller than that of the electron or positron inside the fermionic loop.
+Numerical tests show that in the whole range of densities and temperatures studied here, we effectively have $\omega\lesssim q\ll p\simeq\epsilon$; therefore, the hard-thermal-loop approximation to the polarization tensor works very well.
+In the regime of low-frequency scattering $(x=\omega/q\ll 1)$, relevant for electron-ion and positron-ion collisions, the longitudinal and transverse components (assuming ultrarelativistic electrons and positrons) reduce to
+$\Pi_{L}(q,\omega)=q_{D}^{2}\chi_{l},\quad\Pi_{T}(q,\omega)=q_{D}^{2}\chi_{t},$ (72)
 with ${\rm Re}\chi_{l}=1-x^{2}$, ${\rm Im}\chi_{l}=-{\pi x}/{2}$, ${\rm Re}\chi_{t}=x^{2}$, and ${\rm Im}\chi_{t}={\pi x}/{4}.$ Here, $q_{D}$ is the Debye wave number, given by
-$q_{D}^{2}=-\frac{4e^{2}}{\pi}\int_{0}^{\infty}d\epsilon\,p\,\epsilon\left(\frac{\partial f^{0+}}{\partial\epsilon}+\frac{\partial f^{0-}}{\partial\epsilon}\right).$ (73)
-III.3 The collision rates
-In this section, we summarize the results of computations of the collision rates given by Eqs.
+$q_{D}^{2}=-\frac{4e^{2}}{\pi}\int_{m}^{\infty}d\epsilon\,p\,\epsilon\left(\frac{\partial f^{0+}}{\partial\epsilon}+\frac{\partial f^{0-}}{\partial\epsilon}\right).$ (73)
+III.3 Collision rates
+In this section, we summarize the results of the computations of the collision rates given by Eqs.
 (39)–(42).
@@
 (32) of Ref.
-Harutyunyan2016 , which corresponds to the limit of static scattering.
+Harutyunyan2016 , which corresponds to the limit of static scattering (this approximation is well justified for ions with $Z_{i}>1$).
 The final formulas read
 $\nu_{pi/ei}$ $=$ $\frac{e^{4}Z_{i}^{2}n_{i}}{2\pi l^{\pm}}\!\int_{m}^{\infty}\!d\epsilon\,f^{0\pm}(\epsilon)\left[1-f^{0\pm}(\epsilon)\right]$ (74)
-$\int_{0}^{2p}dq\,q^{3}S_{i}(q)F_{i}^{2}(q)\frac{4\epsilon^{2}-q^{2}}{|q^{2}+\Pi_{L}|^{2}}.$
+$\times\int_{0}^{2p}dq\,q^{3}S_{i}(q)F_{i}^{2}(q)\frac{4\epsilon^{2}-q^{2}}{|q^{2}+\Pi_{L}|^{2}}.\quad$
 The electron-positron collision rate is computed in Appendix B; the final result reads
@@
 where the first and second terms represent the scattering and annihilation diagrams, respectively.
-Note that the variable $\epsilon$ stands for the energy of the initial state electron in both integrals, whereas the variable $\epsilon^{\prime}$ represents the energy of the final electron in the scattering integral and that of the initial positron in the annihilation integral.
+Note that the variable $\epsilon$ stands for the energy of the initial state electron in both integrals, whereas the variable $\epsilon^{\prime}$ represents the energy of the initial positron in the scattering integral and that of the final electron in the annihilation integral.
 Thus, the variables $q$ and $\omega$ are the momentum and energy of the transferred plasmon in both cases, but the integration limits are different.
@@
 $\tilde{q}_{\pm}={q}_{\pm},\quad\tilde{q}^{\prime}_{\pm}=\left|\sqrt{(\epsilon^{\prime}-{\omega})^{2}-m^{2}}\pm\sqrt{\epsilon^{\prime 2}-m^{2}}\right|.$ (85)
-The lower bound for $\omega$ in the annihilation integral is $\omega_{\rm min}={\rm max}\{\epsilon,\epsilon^{\prime\prime}\}+m$, since the plasmon produced in the annihilation carries the total energy of the electron–positron pair.
-Because the matrix element for $pe$ collisions is the same as the one for epep collisions, see Appendix A, we see from Eqs.
-(41) and (42) that the $pe$ collision rate can be obtained from the $ep$ collision rate by simply replacing $f^{0+}\leftrightarrow f^{0-}$ and $l^{-}\rightarrow l^{+}$, therefore
+The lower bound for $\omega$ in the annihilation integral is $\omega_{\rm min}={\rm max}\{\epsilon,\epsilon^{\prime}\}+m$ since the plasmon produced in the annihilation carries the total energy of the electron–positron pair.
+Because the matrix element for $pe$ collisions is the same as the one for $ep$ collisions (see Appendix A), we see from Eqs.
+(41) and (42) that the $pe$ collision rate can be obtained from the $ep$ collision rate by simply replacing $f^{0+}\leftrightarrow f^{0-}$ and $l^{-}\rightarrow l^{+}$; therefore,
 $\nu_{pe}$ $=$ $\frac{e^{4}}{(2\pi)^{3}l^{+}}\!\int_{m}^{\infty}\!d\epsilon\!\int_{m}^{\infty}\!d\epsilon^{\prime}\!\int_{m-\epsilon^{\prime}}^{\epsilon-m}d\omega\,f^{0+}(\epsilon)f^{0-}(\epsilon^{\prime})$ (86)
@@
 IV Numerical results
-We performed numerical calculations of the electrical conductivities from Eq.
-(49) in the regime relevant to neutron star outer crusts, considering a plasma composed of a single ion species, $^{\makebox[7.97226pt][r]{$\scriptstyle 12$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{C}$}, across the entire temperature–density range studied.
-Below, we first present the phase structure of crustal matter, followed by a discussion of our new conductivity results.
-For practical astrophysical applications, all numerical values are given in c.g.s.
-units.
+We performed numerical calculations of the relaxation times using Eqs.
+(33), (34) and (39)–(42), and the electrical conductivities from Eq.
+(49) in the regime relevant to neutron star outer crusts, considering a plasma composed of a single ion species across the entire temperature–density range studied.
+Below, we first present the phase structure of crustal matter, followed by a discussion of our new results for relaxation times and conductivities.
+A more detailed discussion including the phase structure of the crust and relaxation times is provided for $^{\makebox[7.97226pt][r]{$\scriptstyle 12$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{C}$} nuclei, but the final results for conductivities are also presented for $^{\makebox[7.97226pt][r]{$\scriptstyle 56$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Fe}$} and $^{\makebox[3.98613pt][r]{$\scriptstyle 4$}}_{\makebox[3.98613pt][r]{$\scriptstyle$}}\mathrm{He}$} nuclei.
+For practical astrophysical applications, all numerical values are given in in centimeter-gram-second units.
 IV.1 Phase diagram of crustal plasma
-Figure 3: The temperature-density phase diagram of crustal plasma composed of carbon.
-Ionic component of plasma forms a Boltzmann gas above the Coulomb temperature $T_{C}$, a classical liquid at $T_{p}\leq T\leq T_{C}$, and a quantum liquid below the plasma temperature $T_{p}$.
+Figure 3: Temperature-density phase diagram of crustal plasma composed of carbon.
+The ionic component of plasma forms a Boltzmann gas above the Coulomb temperature $T_{C}$, a classical liquid at $T_{p}\leq T\leq T_{C}$, and a quantum liquid below the plasma temperature $T_{p}$.
 Electrons become degenerate below the Fermi temperature $T_{F}$.
 The three curves around $T_{F}$ correspond to temperatures where the ratio of the total electron density $n^{-}$ to the net electron density $n_{e}$ reaches the value $r$.
+Figure 4: Electron chemical potential $\mu_{e}$ (panel a) and the Debye wave number $q_{D}$ (panel b) as functions of the temperature for three values of the density indicated in the plot.
+The dotted lines show the same quantities in the case where positrons are neglected.
 To describe the state of the crustal plasma, we introduce the following physical quantities.
-The ion mass is defined as $M=Am_{n}$, where $A$ is the ion mass number, and $m_{n}\simeq 939$ MeV is the average nucleon mass.
-The ion number density $n_{i}$ is related to the density of matter $\rho$ by $n_{i}=\rho/M$, and we also define the radius of the spherical volume per ion (so-called the Wigner Seitz cell Shapiro1986 ) as $a=(4\pi n_{i}/3)^{-1/3}$.
+The ion mass is defined as $M=Am_{n}$, where $A$ is the ion mass number and $m_{n}\simeq 939$ MeV is the average nucleon mass.
+The ion number density $n_{i}$ is related to the density of matter $\rho$ by $n_{i}=\rho/M$, and we also define the radius of the spherical volume per ion (the so-called Wigner Seitz cell Shapiro1983 ) as $a=(4\pi n_{i}/3)^{-1/3}$.
 The state of ions with charge number $Z$ is determined by the value of the Coulomb plasma parameter
@@
 The present study focuses on the portion of the phase diagram above the plasma temperature.
+Figure 5: Dependence of electron (solid lines) and positron (dashed lines) relaxation times on density for the temperature fixed at (a) $T=1$ MeV; and (b) $T=10$ MeV.
+The dotted lines show the electron-ion relaxation time $\tau_{ei}=\epsilon/\nu_{ei}$, the dash-dotted lines show the electron-positron relaxation time $\tau_{ep}=\epsilon/2\nu_{ep}$, and the dash–double-dotted lines show the positron-electron relaxation time $\tau_{pe}=\epsilon/2\nu_{pe}$.
 Electrons become degenerate below the Fermi temperature $T_{F}=\epsilon_{F}-m$, where the Fermi energy is defined as $\epsilon_{F}=(p_{F}^{2}+m^{2})^{1/2}$, and the Fermi momentum is given by $p_{F}=(3\pi^{2}n_{e})^{1/3}$.
-At finite temperatures, once the thermal energy of matter becomes sufficient to allow for creation of electron-positron pairs, a thermal population of positrons appears.
-The electron number excess over positrons $n_{e}$ is now the relevant quantity that determine the charge neutrality condition $n_{e}=Zn_{i}$.
+At finite temperatures, once the thermal energy of matter becomes sufficient to allow for the creation of electron-positron pairs, a thermal population of positrons appears.
+The electron number excess over positrons, $n_{e}$ is now the relevant quantity that determines the charge neutrality condition $n_{e}=Zn_{i}$.
 The electron chemical potential is found from the value implied by the charge conservation for electron excess $n_{e}=n^{-}-n^{+}$ for any given values of $\rho$ and $T$.
-The positron/electron proper number densities are given by
+The positron and electron proper number densities are given by
 $n^{\pm}=\int\!\frac{2d\bm{p}}{(2\pi)^{3}}f^{0\pm}(\epsilon)=\frac{1}{\pi^{2}}\!\int_{m}^{\infty}\!d\epsilon\,p\epsilon f^{0\pm}(\epsilon),$ (89)
-where the $f^{0\pm}$ are the Fermi distributions of electron and positrons introduced earlier. Clearly, $n^{-}>n_{e}$ when positrons’ population is not negligible.
+where $f^{0\pm}$ are the Fermi distributions of electrons ()(-) and positrons $(+)$ introduced earlier. Clearly, $n^{-}>n_{e}$ when the population of positrons is not negligible.
 In order to assess the relative importance of positrons in the plasma, we define the ratio $r=n^{-}/n_{e}$.
@@
 The ratio $r$ also increases with decreasing density at fixed $T$, as the pair creation rate is closely related to the ratio $T/T_{F}$.
-The shaded region at the top of the phase diagram corresponds to the regime where the number of created positrons, $n^{+}$, already exceeds the number of electrons in the cold $(T\ll T_{F})$ plasma.
+The shaded region at the top of the phase diagram corresponds to the regime where the number of created positrons (or, equivalently, that of created pairs), $n^{+}$, already exceeds the number of electrons in the cold $(T\ll T_{F})$ plasma.
 At temperatures $T\gg T_{F}$, the number of pairs increases approximately cubically with temperature, as implied by Eq.
 (89), yielding $n^{+}\simeq n^{-}\gg n_{e}$ in this regime.
-IV.2 Electrical conductivity from electron-ion and positron-ion collisions
+We also show the electron chemical potential $\mu_{e}$ and the Debye wave number in Fig.
+4 as functions of temperature for three values of the density.
+The dotted lines correspond to the case in which the positron abundance is neglected.
+In the degenerate limit, $T\ll T_{F}$, the positron contribution vanishes, and we recover $\mu_{e}\simeq\epsilon_{F}$ and $q_{D}^{2}\simeq 4e^{2}p_{F}\epsilon_{F}/\pi$, which depend only on the density, as seen in the figure.
+In the opposite, high-temperature limit $T\gg T_{F}$, the chemical potential tends to zero, $\mu_{e}\to 0$, and electrons and positrons contribute nearly equally to the Debye wave number.
+In this regime, Eq.
+(73) yields the approximation $q_{D}^{2}\simeq 16e^{2}T^{2}/\pi.$ This behavior contrasts with the high-temperature scaling $q_{D}^{2}\simeq 4e^{2}\pi n_{e}/T$ obtained when the positron contribution is neglected.
+Consequently, the presence of positrons leads to significantly larger, density-independent Debye wave numbers and therefore to stronger screening of scattering amplitudes compared to the positron-free case.
+IV.2 Relaxation times
+Figure 6: Temperature dependence of electron (solid lines) and positron (dashed lines) relaxation times for the density fixed at (a) $\rho=10^{6}$ g cm-3 and (b) $\rho=10^{8}$ g cm-3.
+The dotted lines show the electron-ion relaxation time $\tau_{ei}=\epsilon/\nu_{ei}$, the dash-dotted lines show the electron-positron relaxation time $\tau_{ep}=\epsilon/2\nu_{ep}$, and the dash–double-dotted lines show the positron-electron relaxation time $\tau_{pe}=\epsilon/2\nu_{pe}$.
+In this section, we discuss the electron and positron relaxation times arising from electron-ion, positron-ion, and electron-positron scattering processes.
+Let us recall that the electron–positron and positron–electron collision rates given by Eqs.
+(75) and (86) contain two distinct microscopic interaction channels: ordinary scattering and pair annihilation–creation processes.
+The latter is found to contribute significantly less to the total collision rate than the genuine scattering channel.
+Figure 5 shows the density dependence of the relaxation times at fixed temperatures $T=1$ MeV (left panel) and $T=10$ MeV (right panel).
+Since the relaxation times are energy dependent $(\tau^{\pm}=\epsilon X^{\pm}$ with $X^{\pm}=\mathrm{const})$, their evaluation depends on the degeneracy of the electrons.
+To distinguish between the degenerate and nondegenerate regimes, we introduce the characteristic temperature $T^{*}\equiv T_{F}/3$, which serves as the transition scale between the two regimes.
+In degenerate matter, $T\leq T^{*}$, the electron relaxation time $\tau^{-}$ is evaluated at the electron Fermi energy $\epsilon_{F}$, while in the nondegenerate ultrarelativistic regime, $T\geq T^{*}$, it is evaluated at the thermal energy $\bar{\epsilon}=3T$.
+The positron relaxation time $\tau^{+}$, by contrast, is always evaluated at the thermal energy.
+In order to assess the relative importance of the separate scattering processes, we show the partial contributions to the relaxation times due to electron-ion and electron-positron collisions.
+At high densities, where the positrons are irrelevant, the electron relaxation times coincide with those due to electron-ion collisions $\tau_{ei}\equiv\epsilon/\nu_{ei}$.
+In this regime $\tau^{-}$ decreases with density because of increasing electron-ion scattering rates; here, we recover the old results of Ref.
+Harutyunyan2016 .
+In the opposite limit of dilute matter, where $T\gg T_{F}$, electron-positron collisions strongly dominate $ei$ and $pi$ collisions, and $\tau^{-}$ and $\tau^{+}$ are close to their limiting values $\tau_{ep}\simeq\tau_{pe}\equiv\epsilon/2\nu_{ep}$, as follows from Eq.
+(38).
+Here $\tau_{ep}$ ($\tau_{pe}$) tends to a density-independent value below the transition point, but increases (decreases) above that point, as the positron population drops; the net electron density increases when matter enters the degenerate regime.
+The total electron relaxation time $\tau^{-}$ interpolates between these two regimes, approaching a density-independent value $\tau_{ep}$ at small densities and decreasing at higher densities.
+The transition point moves to higher densities for larger values of $T$.
+We remark also that the longitudinal and transverse parts of the scattering amplitudes contribute almost equally to the electron-positron collision rates $\nu_{ep}$ and $\nu_{pe}$ in the regime where these are relevant.
+The behavior of the positron relaxation time $\tau^{+}$ is very similar to that of $\tau^{-}$ but with slightly lower values.
+The partial contribution $\tau_{pi}$ is very close to $\tau_{ei}$ in the regime where it is relevant; therefore, it is not shown in the figures.
+Indeed, as indicated by Eqs.
+(43) and (74), the electron-ion and positron-ion collision rates are essentially identical in the regime where positrons are sufficiently abundant, since the integrals over combinations of Fermi functions in the numerators and denominators largely cancel.
+Figure 6 shows the temperature dependence of the relaxation times for several fixed values of the density.
+We see again that, in the strongly degenerate regime $T\ll T_{F}$, we have $\tau^{-}\simeq\tau_{ei}$, which decreases with the temperature due to the ion structure factor $S(q)$ Harutyunyan2016 .
+At intermediate temperatures $T\sim T^{*}$, where the matter is semidegenerate, $\tau_{ei}$ changes its behavior from a decreasing to an increasing function.
+This transition reflects the approximate scaling $\tau_{ei}\propto\epsilon\propto T$.
+In the semidegenerate regime, there exists a temperature window in which electron–positron collisions remain subdominant, i.e., $\tau_{ei}\gtrsim\tau_{ep}$.
+As a result, the effective electron relaxation time $\tau^{-}$ increases in this domain.
+However, at higher temperatures $T\gg T_{F}$, the electron-positron collisions become dominant, and we have $\tau^{-}\simeq\tau_{ep}\propto T^{-1}$, which follows from dimensional arguments.
+Thus, as a result of nontrivial convolution of electron-ion and electron-positron collisions, the electron relaxation time $\tau^{-}$ shows a nonmonotonic behavior with temperature, being a decreasing function in both limiting cases (with a higher slope in the nondegenerate regime), and increasing in the transitional regime which interpolates between these two limiting cases.
+The behavior of the positron relaxation time $\tau^{+}$ is very similar to that of $\tau^{-}$ in the transitional and nondegenerate regimes where it is relevant, albeit with slightly smaller values.
+IV.3 Electrical conductivity from electron-ion and positron-ion collisions
+Figure 7: Dependence of electronic (dashed lines) and positronic (dash-double-dotted lines) partial conductivities and their sum (solid lines) for three values of the temperature indicated in the plot.
+These are the conductivities resulting only from eiei and pipi collisions.
+The dotted lines show the conductivity from Ref.
+Harutyunyan2016 where positron contribution is neglected.
+Figure 8: Temperature dependence of electronic (dashed lines) and positronic (dash-double-dotted lines) partial conductivities and their sum (solid lines) for three values of the density indicated in the plot.
+These are the conductivities resulting only from eiei and pipi collisions.
+The dotted lines show the conductivity from Ref.
+Harutyunyan2016 where positron contribution is neglected.
 For clarity, we first consider the conductivities that include only electron–ion and positron–ion collisions.
 The full conductivities, which also account for electron–positron collisions, will be addressed in the next subsection.
-As indicated by Eqs.
-(43) and (74), the relaxation times for electrons and positrons are essentially identical in the regime where positrons are sufficiently abundant, since the integrals over combintations of Fermi functions in the numerators and denominators largely cancel.
-Consequently, the primary distinction between the electron and positron contributions to the electrical conductivity arises from the Fermi distribution functions in Eq.
+As discussed above, $\tau_{ei}\simeq\tau_{pi}$ in the regime where positrons are sufficiently abundant; therefore, the primary distinction between the electron and positron contributions to the electrical conductivity arises from the Fermi distribution functions in Eq.
 (49).
@@
 where $\tau_{F}\equiv\tau^{-}(\epsilon_{F})$. This is the well-known Drude formula for the electrical conductivity in the theory of metals Abrikosov:Fundamentals .
-In the opposite limit of high temperatures $T\gg T_{F}$, we have also $T\gg\mu$, electrons and positrons are ultrarelativistic with the average thermal energy $\bar{\epsilon}\simeq 3T$, and their number densities are almost equal $n^{+}\simeq n^{-}\equiv\bar{n}\sim\bar{\epsilon}^{3}$.
-In this case $f^{0+}\simeq f^{0-}\equiv f^{0}\ll 1$, and Eq.
+In the opposite limit of high temperatures $T\gg T_{F}$, we also have $T\gg\mu_{e}$, electrons and positrons are ultrarelativistic with the average thermal energy $\bar{\epsilon}\simeq 3T$, and their number densities are almost equal, $n^{+}\simeq n^{-}\equiv\bar{n}\sim\bar{\epsilon}^{3}$.
+In this case, $f^{0+}\simeq f^{0-}\equiv f^{0}\ll 1$, and Eq.
 (49) can be approximated as
@@
 $\bar{n}\bar{F}=\frac{1}{\pi^{2}}\int_{0}^{\infty}{p^{2}dp}\,F(\epsilon)f^{0}(\epsilon).$ (92)
-Thus, the formulas for conductivity in both – strongly degenerate and nondegenerate regimes have the same form, but contain different energy scales, specifically $\epsilon_{F}$ in the degenerate regime and $\bar{\epsilon}\simeq 3T$ in the nondegenerate, ultrarelativistic regime.
+Thus, the formulas for conductivity in both strongly degenerate and nondegenerate regimes have the same form but contain different energy scales, specifically, $\epsilon_{F}$ in the degenerate regime and $\bar{\epsilon}\simeq 3T$ in the nondegenerate, ultrarelativistic regime.
 The relaxation time (74) in the physical regimes of interest was discussed in detail in Ref.
 Harutyunyan2016 .
-On average, it was found that for carbon, $\tau$ varies with density and temperature according to the scaling $\tau(\epsilon)\propto\epsilon^{2}\rho^{-0.9}T^{-0.2}$.
+On average, it was found that, for carbon, $\tau$ varies with density and temperature according to the scaling $\tau(\epsilon)\propto\epsilon^{2}\rho^{-0.9}T^{-0.2}$.
 Using this scaling and the rough estimates $n_{e}\sim\epsilon_{F}^{3}$ and $\bar{n}\sim\bar{\epsilon}^{3}$, Eqs.
@@
 This corresponds approximately to $\sigma\propto\rho^{0.5}T^{-0.2}$ in the degenerate regime and $\sigma\propto\rho^{-1}T^{4}$ in the nondegenerate regime.
-Figure 4: Dependence of electronic (dashed lines) and positronic (dash-double-dotted lines) partial conductivities and their sum (solid lines) for three values of the temperature indicated in the plot.
-These are the conductivities resulting only from eiei and pipi collisions.
-The dotted lines show the conductivity from Ref.
-Harutyunyan2016 where positrons’ abundance in the matter is neglected.
-Figure 5: The temperature dependence of electronic (dashed lines) and positronic (dash-double-dotted lines) partial conductivities and their sum (solid lines) for three values of the density indicated in the plot.
-These are the conductivities resulting only from eiei and pipi collisions.
-The dotted lines show the conductivity from Ref.
-Harutyunyan2016 where positrons’ abundance in the matter is neglected.
 Figure 7 shows the partial conductivities $\sigma^{-}$ and $\sigma^{+}$, as well as the total conductivity $\sigma=\sigma^{-}+\sigma^{+}$ as functions of density for three temperature values, which are chosen to cover the range from the degenerate regime ($T=0.1$ MeV) to the nondegenerate regime ($T=10$ MeV).
-The intermediate value $T=1$ MeV is a representative of the transition between these regimes, which occurs around $\rho\simeq 10^{8}$ g cm-3, see Fig.
+The intermediate value $T=1$ MeV is a representative of the transition between these regimes, which occurs around $\rho\simeq 10^{8}$ g cm-3; see Fig.
 3.
 For comparison, we also plot the conductivities obtained previously in Ref.
-Harutyunyan2016 , where the presence of positrons in the matter was not taken into account.
+Harutyunyan2016 , where the presence of positrons in matter was not taken into account.
+Figure 9: Same as Fig. 7, but with the inclusion of epep collisions.
+Figure 10: Same as Fig. 8, but with the inclusion of epep collisions.
 The qualitative behavior of the conductivity in the degenerate regime is opposite to that in the transitional and nondegenerate regimes.
-This contrast can be understood by noting that at a temperature of $T=0.1$ MeV, positrons are effectively absent, since the temperature is far below the electron rest mass.
+This contrast can be understood by noting that at a temperature of $T=0.1$ MeV, positrons are effectively absent since the temperature is far below the electron rest mass.
 The positron distribution is also strongly suppressed at $\rho\geq 10^{8}\,\mathrm{g\,cm^{-3}}$ for $T=1$ MeV due to Pauli blocking: As the Fermi energy increases, the number of kinematically accessible states for pair creation diminishes.
@@
 In this regime, the behavior of the conductivity reverses: At fixed temperature, it decreases with density, following the power law $\sigma\propto\tau\propto\rho^{-1}$.
-This occurs because the total number of electron–positron pairs, $\bar{n}$, is determined by temperature alone and is independent of the matter density.
+This occurs because in this regime, the total number of electron–positron pairs, $\bar{n}$, is determined by temperature alone and is independent of the matter density.
 The two conductivity branches representing the degenerate and nondegenerate regimes merge smoothly at a minimum located at their intersection.
-This transition point corresponds approximately to the temperature $T\simeq T_{F}/3$, where the two characteristic energy scales become comparable.
+This transition point corresponds approximately to the temperature $T\simeq T^{*}$, where the two characteristic energy scales become comparable.
 Let us now turn to the temperature dependence of the conductivity (see Fig.
-5).
+8).
 The low-temperature branches again reproduce the results of Ref.
@@
 In Ref.
-Harutyunyan2016 , the conductivity was found to reach a minimum at the transition point $T\simeq T_{F}/3$, resulting from the reversed temperature dependence of the relaxation time in the nondegenerate regime.
+Harutyunyan2016 , the conductivity was found to reach a minimum at the transition point $T\simeq T^{*}$, resulting from the reversed temperature dependence of the relaxation time in the nondegenerate regime.
 As seen from Fig.
-5, the position of the minimum around $T\simeq T_{F}/3$ remains unchanged also when positrons are present.
+8, the position of the minimum around $T\simeq T^{*}$ also remains unchanged when positrons are present.
 However, the slope of the curves above the transition point rises abruptly, which is the consequence of the fast opening of the phase space for the pair creation with increasing temperature, as seen also from the phase diagram in Fig.
 3.
-In contrast to the previous scaling $\sigma\propto T^{0.8}$ in the nondegenerate regime, we find here $\sigma\propto T^{4}$.
+In contrast to the previous scaling $\sigma\propto T^{0.8}$ in the nondegenerate regime, here we find $\sigma\propto T^{4}$.
 To give a quantitative example, we compare the results with and without positrons at the lowest density considered here, $\rho=10^{6}\,\mathrm{g\,cm^{-3}}$.
-The inclusion of positrons increases the conductivity rougly by two orders of magnitude at $T=1$ MeV and by six orders of magnitude at $T=10$ MeV.
-IV.3 Electrical conductivity including electron-positron collisions
-Figure 6: Same as Fig. 4, but with inclusion of epep collisions.
+The inclusion of positrons increases the conductivity roughly by 2 orders of magnitude at $T=1$ MeV and by 6 orders of magnitude at $T=10$ MeV.
+IV.4 Electrical conductivity including electron-positron collisions
 We now discuss the electrical conductivities obtained in the full calculation, where electron–positron collisions are included as well.
-Let us recall that the electron–positron and positron–electron collision rates given by Eqs.
-(75) and (86) contain two distinct microscopic interaction channels: ordinary scattering and pair annihilation and creation processes.
-The latter is found to contribute significantly less to the total collision rate than the pure scattering channel.
 Figure 9 shows the partial conductivities $\sigma^{-}$ and $\sigma^{+}$ and their sum $\sigma$ as functions of density for the same temperature values as in Fig.
-4.
-Comparing these two figures, we see that the effect of electron-positron collisions becomes crucial in the low-density and high-temperature regime of non-degenerate matter where $T\gg T_{F}$, as expected.
-Here, the conductivity becomes density-independent.
-This behavior can be explained by the fact that $ep$ collisions heavily outweigh $ei$ and $pi$ collisions in this regime.
-As can be seen from Eq.
-(38), in this case the relaxation times contain only the collision rates $\nu_{ep}\simeq\nu_{pe}$, which are density independent in the regime $\mu_{e}\ll T$, thus resulting in density-independent conductivities in accordance with Eq.
-(91).
+7.
+Comparing these two figures, we see that the effect of electron-positron collisions becomes crucial in the low-density and high-temperature regime of nondegenerate matter where $T\gg T_{F}$, as expected.
+Here, the conductivity becomes density independent, as a consequence of density independence of relaxation times in this regime and the relation (91).
 In the degenerate regime, the electron-positron collisions are suppressed, and we recover the results of the previous section.
-Although the location of the minimum in the total conductivity remains unaffected by electron-positron collisions, these are non-negligible already at $T\geq T_{F}/3$.
-Figure 7: Same as Fig. 5, but with inclusion of epep collisions.
+Although the location of the minimum in the total conductivity remains unaffected by electron-positron collisions, these are non-negligible already at $T\geq T^{*}$.
+Figure 11: Total conductivity as a function of density for three different ionic compositions and two values of the temperature indicated in the plot.
+Figure 12: Total conductivity as a function of temperature for three different ionic compositions and three values of the density indicated in the plot.
 Turning to the temperature dependence of conductivities shown in Fig.
-7, we see again that the main difference as compared to Fig.
-5 appears in the high temperature regime above the minimum point $T\simeq T_{F}/3$.
+10, we see again that the main difference as compared to Fig.
+8 appears in the high-temperature regime above the minimum point $T\simeq T^{*}$.
 At very high temperatures, where $T\gg T_{F}$, the conductivity $\sigma$ becomes essentially independent of density.
 On dimensional grounds, one then finds the scaling $\sigma\propto T$ in this regime, which agrees with the high-$T$ asymptotics shown in Fig.
-7.
+10.
 This behavior is particularly evident at $\rho=10^{6}~\mathrm{g\,cm^{-3}}$.
-Immediately above the transition temperature $T\simeq T_{F}/3$, $\sigma$ follows the power-law dependence $\sigma\propto T^{4}$ as the collisions with ions are still dominant here.
+Immediately above the transition temperature $T\simeq T^{*}$, $\sigma$ follows the power-law dependence $\sigma\propto T^{4}$, as the collisions with ions are still dominant here.
 However, this fast increase is mitigated at higher temperatures as a result of intense electron-positron collisions, which surpass the $ei$ and $pi$ collision rates.
-As a numerical example, we find that at $\rho=10^{6}$ g cm-3 the ratios of conductivities we find to those computed earlier in Ref.
-Harutyunyan2016 are factor $20$ and $30$ for $T=1$ and $T=10$ MeV, respectively.
+As a numerical example, at $\rho=10^{6}$ g cm-3, the ratios of conductivities we find to those computed earlier in Ref.
+Harutyunyan2016 are factors of $20$ and $30$ for $T=1$ and $T=10$ MeV, respectively.
 These factors decrease at higher densities but still reach nearly an order of magnitude across most of the outer crust at $T=10$ MeV.
-Concluding this section, we emphasize that the accounting for creation of positrons is essential for accurate determination of the transport properties in the outer crust of neutron stars, already at temperatures close to the Fermi temperature.
+To give an estimate of how our results would change depending on the crust composition, we show the density and temperature dependence of the total conductivity $\sigma$ for a crust composed of $^{\makebox[3.98613pt][r]{$\scriptstyle 4$}}_{\makebox[3.98613pt][r]{$\scriptstyle$}}\mathrm{He}$}, $^{\makebox[7.97226pt][r]{$\scriptstyle 12$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{C}$} and $^{\makebox[7.97226pt][r]{$\scriptstyle 26$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Fe}$} in Figs.
+11 and 12, respectively.
+The qualitative behavior of the conductivity is similar for these three elements.
+In all cases, there is a minimum (both in the density and temperature dependence) at the point where $T\simeq T^{*}$.
+At densities above or temperatures below this point, the conductivity is dominated by electron-ion collisions and scales as $\sigma\propto Z^{-1}$, as follows from Eqs.
+(74) and (90) (recall that $n_{e}=Zn_{i}=Z\rho/Am_{n}$ is essentially independent of $Z$).
+In the high-temperature and low-density regime, the conductivity becomes independent of both density and nuclei species, approaching a temperature-dependent value $\sigma\,[{\rm s}]\simeq 3\times 10^{22}\,T\,[{\rm MeV}]$.
+Thus, the composition dependence of conductivity is less pronounced in the high-temperature regime where electron-positron collisions play the main role in relaxation processes.
+Concluding this section, we emphasize that accounting for creation of positrons is essential for the accurate determination of the transport properties in the outer crust of neutron stars, already at temperatures close to the Fermi temperature.
 V Summary
@@
 The novelty of our work is the inclusion of the contribution of positrons to the conductivity, which becomes increasingly important as the temperature increases.
-In this regime, the conductivity is dominated by thermally created electron-positron pairs which scatter off the correlated nuclei as well as off each other via screened electromagnetic force.
+In this regime, the conductivity is dominated by thermally created electron-positron pairs, which scatter off the correlated nuclei as well as off each other via screened electromagnetic force.
 Solving a system of coupled Boltzmann kinetic equations for electrons and positrons, we expressed the conductivity in terms of electron and positron relaxation times.
-These contain electron-ion, positron-ion and electron-positron collisions rates, which were computed by using the relevant leading-order scattering diagrams which include all screening factors of strongly correlated crustal plasma.
-Numerical results were obtained for conductivities for crustal matter composed of $^{\makebox[7.97226pt][r]{$\scriptstyle 12$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{C}$}.
+These contain electron-ion, positron-ion and electron-positron collision rates, which were computed by using the relevant leading-order scattering diagrams, which include all screening factors of strongly correlated crustal plasma.
+Numerical results were obtained for conductivities mainly for crustal matter composed of $^{\makebox[7.97226pt][r]{$\scriptstyle 12$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{C}$}.
 We found that the conductivity increases with density in the degenerate regime, in agreement with earlier results of Ref.
-Harutyunyan2016 , but in the nondegenerate regime it shows a decreasing behavior, in contrast to the results of previous studies.
+Harutyunyan2016 , but in the nondegenerate regime, it shows a decreasing behavior, in contrast to the results of previous studies.
 As a result, the conductivity crosses a minimum at the transition point from the nondegenerate to the degenerate regime.
-As a function of temperature, the conductivity grows with a power-law $\sigma\propto T^{4}$ in the semi-degenerate regime close to the transition temperature $T\simeq T_{F}/3$, because of intense creation of thermal electron-positron pairs.
+As a function of temperature, the conductivity grows with a power law $\sigma\propto T^{4}$ in the semidegenerate regime close to the transition temperature $T\simeq T_{F}/3$ because of intense creation of thermal electron-positron pairs.
 In this region, the number of pairs increases as $\propto T^{3}$, but the scattering is still mainly due to ions.
@@
 We find that the inclusion of positrons enhances the electrical conductivity in the hot regions of the outer crust by factors of about 10.
-The results obtained in this work therefore, underscore the importance of accounting for positrons in the transport properties of heated plasma in neutron-star crusts.
-Acknowledgements
+The results obtained in this work therefore underscore the importance of accounting for positrons in the transport properties of heated plasma in neutron-star crusts.
+The present study can be extended by considering more realistic, temperature-dependent crustal compositions that include multiple nuclear species in statistical equilibrium, as well as light nuclear clusters.
+Such an approach would provide a more accurate description of the microphysics in neutron star crusts, where the distribution of nuclei can vary significantly with density and temperature.
+In addition, the formalism developed here may be applied to improve calculations of various transport coefficients, including thermal conductivity, viscosities, and thermoelectric coefficients, in hot neutron star matter.
+Incorporating the effects of positrons in these calculations is particularly important, as they can significantly modify screening and collision rates, thereby influencing the transport properties of the warm crust.
+These extensions would allow for a more comprehensive and realistic modeling of the thermal and transport behavior of the hot transient state of compact stars at subnuclear densities.
+ACKNOWLEDGMENTS
 The authors acknowledge support from the Collaborative Research Grant No.
@@
 2023/51/B/ST9/02798.
-Appendix A The electron-positron scattering amplitude
-In this Appendix, we will provide the details of the calculation of the spin-averaged electron-positron scattering amplitude. From Eqs. (50)–(III.1) we have
+Appendix A Electron-positron scattering amplitude
+In this appendix, we will provide the details of the calculation of the spin-averaged electron-positron scattering amplitude. From Eqs. (50)–(III.1), we have
 $\big|{\cal M}_{12\to 34}^{ep}\big|^{2}$ $=$ $\big|{\cal M}_{12\to 34}^{ep,t}\big|^{2}+\big|{\cal M}_{12\to 34}^{ep,s}\big|^{2}-2{\rm Re}\left({\cal M}_{12\to 34}^{ep,t}{\cal M}_{12\to 34}^{ep,s*}\right),$ (93)
@@
 The spin-averaged current-current couplings are given by
-$\frac{1}{2}\sum\limits_{s_{1}s_{3}}J_{\mu}J_{\nu}^{*}$ $=$ $\frac{e^{*2}}{2\epsilon_{1}\epsilon_{3}}\Big[p_{1\mu}p_{3\nu}+p_{1\nu}p_{3\mu}-g_{\mu\nu}(\epsilon_{1}\epsilon_{3}-{\bm{p}}_{1}\cdot{\bm{p}}_{3}-m^{2})\Big].$ (101)
+$\frac{1}{2}\sum\limits_{s_{1}s_{3}}J_{\mu}J_{\nu}^{*}$ $=$ $\frac{e^{*2}}{2\epsilon_{1}\epsilon_{3}}\Big[p_{1\mu}p_{3\nu}+p_{1\nu}p_{3\mu}-g_{\mu\nu}(\epsilon_{1}\epsilon_{3}-{\bm{p}}_{1}\cdot{\bm{p}}_{3}-m^{2})\Big],$ (101)
 $\frac{1}{2}\sum\limits_{s_{2}s_{4}}J^{\prime}_{\mu}J^{\prime*}_{\nu}$ $=$ $\frac{e^{*2}}{2\epsilon_{2}\epsilon_{4}}\Big[p_{2\mu}p_{4\nu}+p_{2\nu}p_{4\mu}-g_{\mu\nu}(\epsilon_{2}\epsilon_{4}-{\bm{p}}_{2}\cdot{\bm{p}}_{4}-m^{2})\Big].$ (102)
-Substituting the relevant component of the currents into Eq. (97) we obtain for the $t$-channel terms
+Substituting the relevant components of the currents into Eq. (97), we obtain for the $t$-channel terms
 $\frac{1}{4}\sum\limits_{\rm spins}|{\cal M}^{t}_{L}|^{2}$ $=$ $\frac{e^{*4}}{4\epsilon_{1}\epsilon_{2}\epsilon_{3}\epsilon_{4}}\frac{(\epsilon_{1}\epsilon_{3}+{\bm{p}}_{1}\cdot{\bm{p}}_{3}+m^{2})(\epsilon_{2}\epsilon_{4}+{\bm{p}}_{2}\cdot{\bm{p}}_{4}+m^{2})}{|t_{0}|^{2}},$ (103)
@@
 (104) and (105) we used the fact that $\delta^{\perp}_{ik}=\delta_{ik}-{q}_{i}{q}_{k}/{\bm{q}}^{2}$, and ${\bm{p}}_{3}^{\perp}={\bm{p}}_{1}^{\perp}\equiv{\bm{p}}^{\perp}$, ${\bm{p}}_{4}^{\perp}={\bm{p}}_{2}^{\perp}\equiv{\bm{p}}^{\prime\perp}$, which follow from the relations ${\bm{p}}_{3}={\bm{p}}_{1}-{\bm{q}}$, ${\bm{p}}_{4}={\bm{p}}_{2}+{\bm{q}}$.
-We now modify the nominators using the substitutions ${\bm{p}}_{3}={\bm{p}}-{\bm{q}}$, ${\bm{p}}_{4}={\bm{p}}^{\prime}+{\bm{q}}$ and $\epsilon_{3}=\epsilon-\omega$, $\epsilon_{4}=\epsilon^{\prime}+\omega$ (with $p\equiv p_{1}$ and $p^{\prime}\equiv p_{2}$)
+We now modify the numerators using the substitutions ${\bm{p}}_{3}={\bm{p}}-{\bm{q}}$, ${\bm{p}}_{4}={\bm{p}}^{\prime}+{\bm{q}}$ and $\epsilon_{3}=\epsilon-\omega$, $\epsilon_{4}=\epsilon^{\prime}+\omega$ (with $p\equiv p_{1}$ and $p^{\prime}\equiv p_{2}$)
 $\epsilon_{1}\epsilon_{3}+{\bm{p}}_{1}\cdot{\bm{p}}_{3}+m^{2}=2\epsilon^{2}-\epsilon\omega-{\bm{p}}\cdot{\bm{q}},$
@@
 Figure 13: The relative orientation of three vectors $\bm{q}$}, $\bm{p}$} and $\bm{p}$$^{\prime}$ with angles defined by Eq. (107).
-Defining the thee angles between the vectors $\bm{q}$}, $\bm{p}$} and $\bm{p}$$^{\prime}$ (see Fig. 8)
+Defining the three angles between the vectors $\bm{q}$}, $\bm{p}$}, and $\bm{p}$$^{\prime}$ by (see Fig. 13)
 ${\bm{p}}\cdot{\bm{q}}=pq\cos\alpha,\quad{\bm{p}}^{\prime}\cdot{\bm{q}}=p^{\prime}q\cos\theta,\quad{\bm{p}}_{\perp}\cdot{\bm{p}}^{\prime}_{\perp}=p_{\perp}p^{\prime}_{\perp}\cos\varphi,$ (107)
-we obtain for Eq. (106)
+we obtain, for Eq. (106),
 $\frac{1}{4}\sum\limits_{\rm spins}|{\cal M}^{ep,t}_{12\to 34}|^{2}$ $=$ $\frac{e^{*4}}{4\epsilon\epsilon^{\prime}(\epsilon-\omega)(\epsilon^{\prime}+\omega)}\Bigg\{\frac{(2\epsilon^{2}-\epsilon\omega-pq\cos\alpha)(2\epsilon^{\prime 2}+\epsilon^{\prime}\omega+p^{\prime}q\cos\theta)}{|t_{0}|^{2}}$ (108)
@@
 (56), and using the same techniques as above, we obtain
-$\frac{1}{2}\sum\limits_{s_{1}s_{2}}\tilde{J}_{\mu}\tilde{J}_{\nu}^{*}$ $=$ $\frac{e^{*2}}{2\epsilon_{1}\epsilon_{2}}\Big[p_{1\mu}p_{2\nu}+p_{1\nu}p_{2\mu}-g_{\mu\nu}(\epsilon_{1}\epsilon_{2}-{\bm{p}}_{1}\cdot{\bm{p}}_{2}+m^{2})\Big].$ (109)
+$\frac{1}{2}\sum\limits_{s_{1}s_{2}}\tilde{J}_{\mu}\tilde{J}_{\nu}^{*}$ $=$ $\frac{e^{*2}}{2\epsilon_{1}\epsilon_{2}}\Big[p_{1\mu}p_{2\nu}+p_{1\nu}p_{2\mu}-g_{\mu\nu}(\epsilon_{1}\epsilon_{2}-{\bm{p}}_{1}\cdot{\bm{p}}_{2}+m^{2})\Big],$ (109)
 $\frac{1}{2}\sum\limits_{s_{3}s_{4}}\tilde{J}^{\prime}_{\mu}\tilde{J}^{\prime*}_{\nu}$ $=$ $\frac{e^{*2}}{2\epsilon_{3}\epsilon_{4}}\Big[p_{3\mu}p_{4\nu}+p_{3\nu}p_{4\mu}-g_{\mu\nu}(\epsilon_{3}\epsilon_{4}-{\bm{p}}_{3}\cdot{\bm{p}}_{4}+m^{2})\Big],$ (110)
@@
 To check this, it is sufficient to interchange electrons and positrons, which corresponds to the following choice of Dirac 4-currents instead of Eqs.
-(55)–(56)
+(55) and (56),
 $J^{\mu}=-e^{*}\bar{v}_{1}\gamma^{\mu}v_{3},\quad J^{\prime\mu}=-e^{*}\bar{u}_{4}\gamma^{\mu}u_{2},$ (114)
 $\tilde{J}^{\mu}=-e^{*}\bar{v}_{1}\gamma^{\mu}u_{2},\quad\tilde{J}^{\prime\mu}=-e^{*}\bar{u}_{4}\gamma^{\mu}v_{3}.$ (115)
-Starting with these currents, a computation analogous to the one above leads then to exactly the same Eqs.
+Starting with these currents, a computation analogous to the one above leads to exactly the same Eqs.
 (101) and (102), as well as Eqs.
 (109) and (110).
-Therefore, also the final forms of the scattering and annihilation terms in the matrix element are the same.
+Therefore, the final forms of the scattering and annihilation terms in the matrix element are the same.
 Appendix B Evaluating the electron-positron collision rate
-Substituting the electron-positron scattering amplitude (93) in the collision rate (41) we obtain
+Substituting the electron-positron scattering amplitude (93) in the collision rate (41), we obtain
 $\nu_{ep}$ $=$ $\frac{(2\pi)^{-8}}{2l^{-}}\!\int\!d\bm{p}_{1}\,d\bm{p}_{2}\,d\bm{p}_{3}\,d\bm{p}_{4}\,\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4}){\bm{q}}^{2}$ (116)
@@
 (107).
-Similar expression follows from Eq.
+A similar expression follows from Eq.
 (112) for the primed angles.